### v4

# On CNF formulas irredundant with respect to unit clause propagation

Abstract
Abstract
Two CNF formulas are called ucp-equivalent, if they behave in the same way with respect to the unit clause propagation (UCP). A formula is called ucp-irredundant, if removing any clause leads to a formula which is not ucp-equivalent to the original one. As a consequence of known results, the ratio of the size of a ucp-irredundant formula and the size of a smallest ucp-equivalent formula is at most $n^{2}$, where $n$ is the number of the variables. We demonstrate an example of a ucp-irredundant formula for a symmetric definite Horn function which is larger than a smallest ucp-equivalent formula by a factor $\Omega(n/\ln n)$ and, hence, a general upper bound on the above ratio cannot be smaller than this.
1 Introduction
In some contexts, the strength of unit clause propagation (UCP) on the formula is important and formulas representing the same function have to be distinguished according to the behavior of UCP on them. An important example is the set of learned clauses in a SAT solver. Most of these clauses are implicates of the original formula so they do not change the represented function and they are added in order to make UCP stronger. In order to investigate CNF formulas in such a context, we need to compare the strength of UCP on them. In this paper, we consider a corresponding equivalence and irredundancy. Two CNF formulas are called ucp-equivalent, if for every partial assignment $\alpha$, UCP derives the same consistent set of literals from the two formulas together with $\alpha$ or we get a contradiction in both cases. A formula is called ucp-irredundant, if removing any clause makes UCP on the formula weaker and, hence, the obtained formula is not ucp-equivalent to the original one.
UCP on a CNF formula can be precisely described by its dual-rail encoding [1, 2, 3, 4, 5, 6, 7] which represents UCP by a specific Horn formula. In particular, two formulas are ucp-equivalent if and only if their dual-rail encodings are equivalent in the usual sense to represent the same Horn function. It is a well-known result [8] that the ratio of the size of a Horn formula irredundant with respect to representing a given function and the size of a smallest equivalent formula is at most the number of the variables. Horn formulas represent a very narrow class of boolean functions. Using the dual-rail encoding we can transform this result to ucp-irredundant formulas representing an arbitrary boolean function. In this case, we consider the ratio of the size of any ucp-irredundant formula and the size of a smallest ucp-equivalent formula. Since dual-rail encoding changes the number of the clauses, we get a worse bound, namely $2n^{2}$, where $n$ is the number of the variables. A slightly better bound $n^{2}$ can be obtained by a proof used in [9] to prove the same bound for propagation complete formulas since the assumption of propagation completeness is used only to guarantee ucp-equivalence. Moreover, it appears that the factor $n^{2}$ holds both for the number of clauses and the length of the formula.
It is not clear whether a general upper bound on the ratio discussed above better than $n^{2}$ can be obtained. The main result of the paper is an example of a ucp-irredundant formula $\varphi^{\ell}$ of $n$ variables which is larger than the smallest ucp-equivalent formula $\varphi^{*}$ by a factor $\Omega(n/\ln n)$. It follows that a general upper bound on the ratio cannot be smaller than this.
The formulas used in the example represent a symmetric definite Horn function and the construction of the formula $\varphi^{\ell}$ is easy. The hard part of the proof is to derive an upper bound on the size of $\varphi^{*}$ and we show that the size of $\varphi^{*}$ is closely related to the size of a specific covering system. Namely, for $n$ and $k=n/2+O(1)$, we use a system of $(k+1)$-subsets of an $n$-set such that every $k$-set is contained in at least one member of the system. The proof is then finished using a known upper bound on the size of such a system from [10] or a different upper bound from [11, 12]. Although the two bounds are incomparable in general, they are close to each other for $n$ and $k$ as above.
Section 3 defines ucp-equivalence and ucp-irredundancy and Section 4 formulates the results. The property of the formulas needed for the results is reformulated in Section 5 as a property of a hypergraph. The construction of the required hypergraph is described in Section 7 using an auxiliary result from Section 6. The proofs of the results are summarized in Section 8. Conclusion and questions for further research are formulated in Section 9.
2 Preliminaries
2.1 CNF formulas
A literal is a variable or its negation, a clause is a disjunction of literals on different variables, and a CNF formula is a conjunction of clauses. An empty clause denoted $\bot$ is a clause containing no literals and represents a contradiction. A unit clause is a clause containing exactly one literal. If $A$ is a set of variables, let $\mathrm{neg}(A)=\{\neg x\mid x\in A\}$. Since we consider only CNF formulas, we call them formulas for simplicity.
A formula is considered as a set of clauses and a clause is considered as a set of literals. In particular, we use set operations on formulas and clauses. The size of a formula is the number of its clauses. The length of a clause is the number of the literals in it. The length of a formula is the total number of occurences of the literals in it. A clause $C$ is an implicate of a formula $\varphi$, if every assignment of the variables satisfying $\varphi$ satisfies $C$. A clause $C$ is a prime implicate of a formula $\varphi$, if it is an implicate of $\varphi$ and no proper subset of $C$ is an implicate of $\varphi$.
A formula is a Horn formula, if every clause contains at most one positive literal, and a formula is a definite Horn formula, if every clause contains exactly one positive literal.
A partial assignment of the variables is a partial map from the variables to the set $\{0,1\}$ understood as a conjunction of identities of the form $x=a$, where $x$ is a variable and $a\in\{0,1\}$. A partial assignment will be represented by a conjunction of the literals equivalent to these identities. For example, the conjunction of identities $x_{1}=1\wedge x_{2}=0$ is equivalent to the conjunction of literals $x_{1}\wedge\neg x_{2}$.
Unit clause propagation (UCP) is a process of deriving literals or a contradiction implied by the formula using the following process. Assume a formula $\varphi$ and let us initialize a set $U$ as the set of clauses of length at most one contained in $\varphi$. Clauses of length one will be identified with the literals contained in them. All clauses in the initial set $U$ are elements of $\varphi$, so they are clearly consequences of $\varphi$. Assume a clause $C\in\varphi$, and a literal $l\in C$. If $U$ contains the negations of all the literals in $C\setminus\{l\}$, then every satisfying assignment of $\varphi$ satisfies $l$ and, hence, $l$ is a consequence of $\varphi$. Similarly, if $U$ contains the negation of all the literals in $C$, then $\bot$ is a consequence of $\varphi$. Unit clause propagation on a formula $\varphi$ is the process of iteratively extending $U$ by literals and possibly $\bot$ derived as consequences of this simple type. The process terminates if no further literal or $\bot$ can be derived or if $\bot\in U$.
Definition 2.1.
For every CNF formula $\varphi$, let $\mathrm{UCP}(\varphi)$ be
  • the set of literals contained in $\varphi$ as unit clauses or derived by unit clause propagation on $\varphi$, if the contradiction is not reached,
[…]
One can show that $\mathrm{UCP}(\varphi)$ does not depend on the order of the derivations performed by the process of UCP. If $\mathrm{UCP}(\varphi)=\{\bot\}$, then $\varphi$ is unsatisfiable. The opposite is not true. For example, the formula $\varphi_{0}=(x_{1}\vee x_{2})\wedge(x_{1}\vee\neg x_{2})\wedge(\neg x_{1}\vee x_{2})\wedge(\neg x_{1}\vee\neg x_{2})$ is unsatisfiable, however, $\mathrm{UCP}(\varphi_{0})=\emptyset$.
2.2 Covering and Turán numbers
A $k$-set or a $k$-subset is a set or a subset of size $k$. We use known bounds on covering numbers $C(n,r,k)$ [10, 13] and the closely related Turán numbers $T(n,\ell,t)$ [12, 11]. Since there are different conventions to define these numbers, we recall the definitions which we use.
Definition 2.2.
For integers $n>r>k>0$, let the covering number $C(n,r,k)$ be the smallest size of a system of $r$-subsets of an $n$-set such that every $k$-subset is contained in at least one set in the system.
Definition 2.3.
For integers $n>\ell>t>0$, let the Turán number $T(n,\ell,t)$ be the smallest size of a system of $t$-subsets of an $n$-set such that every $\ell$-subset contains at least one set in the system.
By taking complements of the sets, we obtain $C(n,r,k)=T(n,n-k,n-r)$. Note that Turán’s theorem implies a precise value of $T(n,\ell,2)$, if we minimize the number of edges to be removed from a complete graph on $n$ vertices in order to eliminate all complete subgraphs on $\ell$ vertices.
Let us recall several known results on covering numbers in the case $r=k+1$ which we use. Clearly,
$C(n,k+1,k)\geq\frac{1}{k+1}{n\choose k}\;.$ (1)
If $k$ is fixed and $n\to\infty$, then by [14], we have
$C(n,k+1,k)=(1+o(1))\frac{1}{k+1}{n\choose k}\;.$
We are mainly interested in cases when $k=n/2+O(1)$ for which only less precise bounds are known. Let $\mu(n,k)$ be such that
$C(n,k+1,k)=\frac{1}{k+1}{n\choose k}\mu(n,k)$ (2)
and, equivalently,
$T(n,t+1,t)=\frac{1}{t+1}{n\choose t}\mu(n,n-t-1)\;.$ (3)
For an upper bound on $\mu(n,k)$, let us recall the following.
Theorem 2.4 (a special case of Theorem 13.4 from [10]).
For every $n>k\geq 2$, we have
$C(n,k+1,k)\leq\frac{1}{k+1}{n\choose k}(1+\ln(k+1))$ (4)
Together with (1), this implies
$1\leq\mu(n,k)\leq 1+\ln(k+1)$ (5)
for every $n>k\geq 2$. An incomparable upper bound on $\mu(n,k)$ follows from the bound [11] on Turán numbers which improves [12] by a factor of $2$. By this bound, we have
$T(n,t+1,t)\leq\frac{(1+o(1))\ln t}{2t}{n\choose t}$ (6)
for $n\to\infty$ and $t\to\infty$. Combining (6) with (3), substituting $k=n-t-1$ and assuming $n-k\to\infty$, we obtain
$\mu(n,k)\leq\frac{1+o(1)}{2}\ln(n-k-1)\;.$ (7)
Bounds (5) and (7) are complementary in the sense that (5) is better, if $k$ is small, while (7) is better, if $k$ is close to $n$. In Section 8 we need also the following.
Lemma 2.5.
For every $n>k>1$, we have
$\mu(n-1,k-1)$$\leq$ $\mu(n,k)$ (8)
$\frac{n-k+1}{n+1}\mu(n,k)$$\leq$ $\mu(n+1,k)$ (9)
Proof.
By [15, 13], we have
$\frac{n}{k+1}C(n-1,k,k-1)\leq C(n,k+1,k)$
implying (8). Assume a system of $(k+1)$-subsets of a set $X$ of $n+1$ points of size $C(n+1,k+1,k)$ and covering all $k$-subsets. For any $x\in X$, one can create a system of $(k+1)$-subsets of $X\setminus\{x\}$ covering all $k$-subsets by replacing each ocurrence of $x$ in a $(k+1)$-tuple in the system by any element of $X\setminus\{x\}$ not contained in the $(k+1)$-tuple. Hence, we have
$C(n,k+1,k)\leq C(n+1,k+1,k)$
implying (9). ∎
3 Ucp-equivalence and ucp-irredundancy
Let us introduce the following equivalence relation on CNF formulas.
Definition 3.1.
Formulas $\varphi_{1}$ and $\varphi_{2}$ are ucp-equivalent if for every partial assignment $\alpha$, we have $\mathrm{UCP}(\varphi_{1}\wedge\alpha)=\mathrm{UCP}(\varphi_{2}\wedge\alpha)$.
Example 3.2.
Formulas $\varphi_{1}=(a\vee c)\wedge(b\vee c)$ and $\varphi_{2}=(a\vee c)\wedge(\neg a\vee b\vee c)$ are equivalent, but not ucp-equivalent, since $\mathrm{UCP}(\varphi_{1}\wedge\neg b)=\{\neg b,c\}$ and $\mathrm{UCP}(\varphi_{2}\wedge\neg b)=\{\neg b\}$.
Example 3.3.
The formula $\varphi_{1}=(\neg a\vee b)\wedge(\neg b\vee c)\wedge(\neg c\vee a)$ and the formula $\varphi_{2}=(\neg a\vee c)\wedge(\neg c\vee b)\wedge(\neg b\vee a)$ are both equivalent and ucp-equivalent. Every positive literal on the variables $a,b,c$ derives all the remaining positive literals by UCP on both $\varphi_{1}$ and $\varphi_{2}$ and the same is true for the negative literals.
This notion appeared implicitly in analysis of propagation complete (PC) formulas [3, 16], so let us recall their definition. A formula $\varphi$ is called propagation complete, if for every partial assignment $\alpha$ of the variables, UCP from $\varphi\wedge\alpha$ derives a contradiction, if $\varphi\wedge\alpha$ is inconsistent, and all literals implied by $\varphi\wedge\alpha$ otherwise. In other words, a formula is propagation complete, if the strength of UCP on it is maximal among all formulas representing the same function. In particular, every two equivalent PC formulas are ucp-equivalent, since UCP on them is controlled by the function they represent. One can also verify that a formula is PC if and only if it is ucp-equivalent to the set of all its (prime) implicates.
Let us consider the notion of an absorbed clause introduced in [3]. A clause $C$ is absorbed by a formula $\varphi$, if for every $l\in C$, UCP from $\varphi\wedge\mathrm{neg}(C\setminus\{l\})$ derives $l$ or a contradiction. A formula is called minimal PC formula in [3] if none of its clauses is absorbed by the remaining ones. This does not guarantee that no smaller equivalent PC formula exists, however, the formula is irredundant with respect to the behavior of UCP. We use a straightforward generalization of this type of irredundancy obtained by removing the requirement that the formula is PC.
Definition 3.4.
A formula $\varphi$ is ucp-irredundant if any of the following equivalent conditions is satisfied:
  • None of the clauses $C\in\varphi$ is absorbed by $\varphi\setminus\{C\}$.
  • For every $C\in\varphi$, the formulas $\varphi$ and $\varphi\setminus\{C\}$ are not ucp-equivalent.
The notion of absorption can naturally be extended to formulas.
Definition 3.5.
A formula $\varphi_{1}$ is absorbed by the formula $\varphi_{2}$, which will be denoted $\varphi_{1}\leq_{ucp}\varphi_{2}$, if every clause of $\varphi_{1}$ is absorbed by $\varphi_{2}$.
One can verify that $\varphi_{1}\leq_{ucp}\varphi_{2}$ if and only if for every partial assignment $\alpha$, we have $\mathrm{UCP}(\varphi_{1}\wedge\alpha)\subseteq\mathrm{UCP}(\varphi_{2}\wedge\alpha)$ or $\mathrm{UCP}(\varphi_{2}\wedge\alpha)=\{\bot\}$. It follows that the relation $\leq_{ucp}$ is a preorder and any two formulas $\varphi_{1}$ and $\varphi_{2}$ are ucp-equivalent if and only if $\varphi_{1}\leq_{ucp}\varphi_{2}$ and $\varphi_{2}\leq_{ucp}\varphi_{1}$. In particular, $\leq_{ucp}$ defines a partial order on the classes of ucp-equivalent formulas and the class of PC formulas is the maximal class in this order. Since, by definition, $\varphi_{1}\leq_{ucp}\varphi_{2}$ can be tested in polynomial time, also ucp-equivalence can be tested in polynomial time.
By the result [9], if $\varphi$ is a ucp-irredundant PC formula, then the ratio of its size and the size of a smallest PC formula equivalent to $\varphi$ is at most $n^{2}$, where $n$ is the number of the variables. The assumption that the two formulas are PC is used only to guarantee that they behave in the same way with respect to UCP or, equivalently, that the two formulas are ucp-equivalent. It follows that the upper bound on the ratio of their sizes can be generalized to non-PC formulas as follows. The statement is the strongest if $\varphi^{*}$ is the smallest formula ucp-equivalent to $\varphi$, however, the proof does not use this assumption.
Proposition 3.6.
If $\varphi$ is a ucp-irredundant formula on $n$ variables and $\varphi^{*}$ is ucp-equivalent to $\varphi$, then
$|\varphi|\leq n^{2}|\varphi^{*}|$ (10)
and
$\|\varphi\|\leq n^{2}\|\varphi^{*}\|$ (11)
Proof.
Every clause of $\varphi^{*}$ is absorbed by $\varphi$. Let $\varphi^{\prime}$ be the set of clauses in $\varphi$ needed for this for all clauses of $\varphi^{*}$. For each pair $l\in C\in\varphi^{*}$, we need at most $n$ clauses of $\varphi$ to derive $l$ or a contradiction from $\varphi\wedge\mathrm{neg}(C\setminus\{l\})$. This implies $|\varphi^{\prime}|\leq n\|\varphi^{*}\|$. Since $\varphi^{\prime}\subseteq\varphi$, we have
$\varphi\leq_{ucp}\varphi^{*}\leq_{ucp}\varphi^{\prime}\leq_{ucp}\varphi$
which implies that the formula $\varphi^{\prime}$ is ucp-equivalent to $\varphi$. Since $\varphi$ is ucp-irredundant, $\varphi^{\prime}$ is equal to $\varphi$. Hence, we have $|\varphi|\leq n\|\varphi^{*}\|$ implying
$|\varphi|\leq n\|\varphi^{*}\|\leq n^{2}|\varphi^{*}|$
and
$\|\varphi\|\leq n|\varphi|\leq n^{2}\|\varphi^{*}\|$
as required. ∎
UCP on a CNF formula can be represented by a Horn formula called dual-rail encoding [1, 2, 3, 4, 5, 6, 7] although this name can refer to different variants of the formula which differ in the way how reaching a contradiction is handled. For definitness, let us specify a formula suitable for our purposes. Assume, $\varphi=C_{1}\wedge\ldots\wedge C_{m}$ is a formula on variables $X=\{x_{1},\ldots,x_{n}\}$. For every literal $l$ on the variables $X$ introduce a variable $\llbracket l\rrbracket$. A total assignment of these variables satisfying $\neg\llbracket x_{i}\rrbracket\vee\neg\llbracket\neg x_{i}\rrbracket$ for every variable $x_{i}\in X$ encodes the set of the satisfied literals under a partial assignment of the variables $X$. Let
$\mathrm{DR}(\varphi)=\mathrm{DR}(C_{1})\wedge\ldots\wedge\mathrm{DR}(C_{m})\wedge\bigwedge_{i=1}^{n}(\neg\llbracket x_{i}\rrbracket\vee\neg\llbracket\neg x_{i}\rrbracket)$
where for a clause $C=l_{1}\vee\ldots\vee l_{k}$,
$\mathrm{DR}(C)=\bigwedge_{i=1}^{k}\left(\bigwedge_{j\not=i}\llbracket\neg l_{j}\rrbracket\to\llbracket l_{i}\rrbracket\right)$
The set of models of $\mathrm{DR}(\varphi)$ is precisely the set of encodings of partial assignments of the variables $X$ which are closed under UCP. Using this, it is easy to prove the following.
Proposition 3.7.
Two formulas $\varphi_{1}$ and $\varphi_{2}$ are ucp-equivalent if and only if the Horn formulas $\mathrm{DR}(\varphi_{1})$ and $\mathrm{DR}(\varphi_{2})$ are equivalent.
4 The results
The main result is the following separation between the size of a ucp-irredundant formula and a smallest ucp-equivalent formula.
Theorem 4.1.
For every $n$ and $k=n/2+O(1)$, there is a ucp-irredundant formula $\varphi^{\ell}$ of $n$ variables and of size
$|\varphi^{\ell}|=(k+1){n-1\choose k}=\Theta(n){n\choose k}$
and such that a smallest ucp-equivalent formula $\varphi^{*}$ has size
$|\varphi^{*}|\leq 6(k+1)C(n,k+1,k)=O(\ln n){n\choose k}$
In particular, $|\varphi^{\ell}|=\Omega(n/\ln n)|\varphi^{*}|$.
For a proof, we investigate the size of ucp-irredundant formulas ucp-equivalent to the following definite Horn formula $\Psi_{n,k}$.
Definition 4.2.
For any integers $n$, $k$, such that $11\leq k\leq n-$ 1, let $\Psi_{n,k}$ be the CNF formula consisting of all clauses on variables $X=\{x_{1},\ldots,x_{n}\}$ containing $k$ negative literals and one positive literal.
Clearly, $|\Psi_{n,k}|=(n-k){n\choose k}$. A boolean function is called symmetric, if it is invariant under any permutation of its variables.
Lemma 4.3.
Formula $\Psi_{n,k}$ represents the symmetric boolean function $f_{n,k}$ of $n$ variables defined by
$f_{n,k}(x_{1},\ldots,x_{n})=1\Leftrightarrow\left(\sum_{i=1}^{n}x_{i}<k\right)\vee\left(\sum_{i=1}^{n}x_{i}=n\right)$
Proof.
By case inspection concerning the number of variables with value 00 and $1$ in a given assignment. ∎
In particular, we prove the following upper and lower bound on the size of a smallest formula ucp-equivalent to $\Psi_{n,k}$.
Theorem 4.4.
For every $11\leq k\leq n-$ 1, the minimum size (number of clauses) of a formula ucp-equivalent to $\Psi_{n,k}$ is at least
$n\,C(n-1,k,k-1)={n\choose k}\mu(n-1,k-1)\geq{n\choose k}$
and if $k=n/2+O(1)$ then it is at most
$6(k+1)C(n,k+1,k)=6\,{n\choose k}\mu(n,k)=O(\ln n){n\choose k}\;.$
The proof of Theorem 4.1 depends on Theorem 4.4 and the proof of both of them is summarized in Section 8 using constructions presented in this and the following sections. The lower and upper bound in Theorem 4.4 are proven to differ at most by a factor $O(\ln n)$, however, it is plausible to assume that they are in fact closer, see the discussion of this in Section 8.
The lower bound in Theorem 4.4 is proven in Section 5. For the proof of the remaining statements, we construct for infinitely many pairs $(n,k)$ ucp-irredundant formulas $\varphi^{\ell}$ (Section 5) and $\varphi^{*}$ (Section 8) ucp-equivalent to $\Psi_{n,k}$, such that $\varphi^{\ell}$ is as large as possible and $\varphi^{*}$ is as small as possible. We obtain $\varphi^{*}$ which meets the upper bound in Theorem 4.4 and the two formulas together satisfy Theorem 4.1. Each of the formulas $\varphi^{\ell}$ and $\varphi^{*}$ will be constructed as a hypergraph formula $\theta(H)$ where $H$ is a suitable hypergraph, see also [17, 18].
Definition 4.5.
For every $(k+1)$-uniform hypergraph $H$ whose vertices are the variables $X$, the hypergraph formula $\theta(H)$ consists of the clauses $\mathrm{neg}(e\setminus\{x\})\cup\{x\}$ for all hyperedges $e\in H$ and all variables $x\in e$.
Clearly, for every $H$, $\theta(H)\subseteq\Psi_{n,k}$ and $|\theta(H)|=(k+1)|H|$. Considering hypergraph formulas for the upper bound in Theorem 4.4 is restrictive, however, this does not seem to be significant. The lower bound is valid for an arbitrary subset of $\Psi_{n,k}$ and the gap between the lower bound and the upper bound obtained by hypergraph formulas is relatively small.
5 A combinatorial characterization
When considering formulas ucp-equivalent to $\Psi_{n,k}$, we consider only prime formulas (containing only prime implicates). Since the formula $\Psi_{n,k}$ is equal to the set of all its prime implicates, these formulas are subsets of $\Psi_{n,k}$. We present two combinatorial characterizations of the subsets of $\Psi_{n,k}$ which are ucp-equivalent to $\Psi_{n,k}$. The first of them is valid for an arbitrary subset of $\Psi_{n,k}$ and is used for a lower bound on the size of an arbitrary formula ucp-equivalent to $\Psi_{n,k}$. The other characterization is valid for subsets of $\Psi_{n,k}$ in the form of a hypergraph formula $\theta(H)$ defined above and is used later for a corresponding upper bound.
Every clause of $\Psi_{n,k}$ contains $k$ negative literals. It follows that UCP on $\Psi_{n,k}$ and a partial assignment setting at most $k-2$ variables to $1$ derives no additional literals. The same is true for every subset of $\Psi_{n,k}$ and, hence, for a characterization of subsets of $\Psi_{n,k}$ which are ucp-equivalent to $\Psi_{n,k}$, it is sufficient to consider partial assignments setting at least $k-1$ variables to $1$ and possibly some variables to 00.
Definition 5.1.
Assume $\varphi\subseteq\Psi_{n,k}$ and a $(k-1)$-set $A\subseteq X$. Let $G^{d}(\varphi,A)$ be the directed graph on $X\setminus A$ defined as
$G^{d}(\varphi,A)=\{(x_{i},x_{j})\mid\mathrm{neg}(A)\cup\{\neg x_{i},x_{j}\}\in\varphi\}\;.$
If a partial assignment sets all variables in $A$ to $1$ and no other variable is set, then UCP derives no additional literals. If also a variable not in $A$ is set, then UCP can be characterized by $G^{d}(\varphi,A)$ as follows.
Lemma 5.2.
Let $\varphi$ and $A$ be as in Definition 5.1. Assume a partial assignment $\alpha$ setting all variables in $A$ to $1$ and setting a variable $x_{\ell}\in X\setminus A$ to $b\in\{0,1\}$. Then $\beta=\mathrm{UCP}(\varphi\wedge\alpha)$ is consistent and we have
  • If $b=0$, then $\beta$ extends $\alpha$ precisely by setting to 00 the variables $x_{i}$ such that $G^{d}(\varphi,A)$ contains a path from $x_{i}$ to $x_{\ell}$.
  • If $b=1$, then $\beta$ extends $\alpha$ at least by setting to $1$ the variables $x_{i}$ such that $G^{d}(\varphi,A)$ contains a path from $x_{\ell}$ to $x_{i}$.
Proof.
By Lemma 4.3, setting all variables in $X\setminus A$ to $b$ is consistent with $\Psi_{n,k}$ and, hence, also with $\varphi$. It follows that $\beta$ is consistent.
Consider setting all variables in $A$ to $1$. With this partial assignment, the clauses containing a positive literal on a variable from $A$ become satisfied and, hence, can be removed. Since $|A|=k-1$, the remaining clauses form a definite Horn formula without unit clauses. In such a formula, setting some of the variables in $X\setminus A$ to 00 can derive a value of a new variable by UCP only using a clause containing one negative literal. Since the derived value of the variable in this literal is 00, the same is true if the propagation is iterated. If $C\in\varphi$, then setting the variables in $A$ to $1$ in $C$ produces a clause with one negative literal only if $C=\mathrm{neg}(A)\cup\{\neg x_{i},x_{j}\}$ for some $i,j$ and the resulting clause is $\{\neg x_{i},x_{j}\}$. Hence, the clauses relevant for propagating assignments of variables in $X\setminus A$ to 00 after setting all variables in $A$ to $1$ form the 2-CNF formula
$\varphi_{A}=\{C\setminus\mathrm{neg}(A)\mid\mathrm{neg}(A)\subseteq C\in\varphi\}\;.$ (12)
The clause $\{\neg x_{i},x_{j}\}$ is the same as $\neg x_{i}\vee x_{j}$ and is equivalent to the implication $x_{i}\to x_{j}$. UCP using this implication derives $x_{i}=0$, if $x_{j}=0$ and the graph $G^{d}(\varphi,A)$ corresponds precisely to implications in $\varphi_{A}$. This proves the statement concerning $b=0$.
The statement concerning $b=1$ is clear, since $\varphi_{A}$ is a subset of the formula obtained from $\varphi$ by setting variables in $A$ to $1$ and the clauses of $\varphi_{A}$ allow propagation of values $1$ along all directed paths in $G^{d}(\varphi,A)$. ∎
Proposition 5.3.
A formula $\varphi\subseteq\Psi_{n,k}$ is ucp-equivalent to $\Psi_{n,k}$ if and only if for every $(k-1)$-set $A\subseteq X$ the graph $G^{d}(\varphi,A)$ is a strongly connected graph on $X\setminus A$.
Proof.
Let $A$ be any $(k-1)$-subset of $X$ and let $\Psi_{n,k,A}$ be obtained from $\Psi_{n,k}$ by setting the variables in $A$ to $1$. Formula $\Psi_{n,k,A}$ contains the clauses $\neg x_{i}\vee x_{j}$ for every $x_{i},x_{j}\in X\setminus A$. It follows that UCP on $\Psi_{n,k,A}$ with a partial assignment setting a variable $x_{j}\in X\setminus A$ to 00 derives the value 00 for all the remaining variables $x_{i}\in X\setminus A$. The resulting assignment is a satisfying assignment of $\Psi_{n,k,A}$, so contradiction is not derived.
Assume, $\varphi\subseteq\Psi_{n,k}$ is ucp-equivalent to $\Psi_{n,k}$. Since UCP on $\Psi_{n,k,A}$ derives value 00 for all variables in $X\setminus A$ if one of them is set to 00, the same should be true for the formula obtained from $\varphi$ by setting the variables in $A$ to $1$. By Lemma 5.2 with $b=0$, this implies that $G^{d}(\varphi,A)$ contains a path in both directions between any two variables in $X\setminus A$, so it is strongly connected.
Assume that $G^{d}(\varphi,A)$ is strongly connected. Let $C\in\Psi_{n,k}\setminus\varphi$ and let us prove that $C$ is absorbed by $\varphi$. Let $l\in C$ and consider a partial assignment $\alpha$ of the variables falsifying all the literals in $C\setminus\{l\}$. We have to show that UCP on $\varphi$ derives $l$ from $\alpha$.
If $l$ is positive, let $g\in C\setminus\{l\}$ be arbitrary. If $l$ is negative, let $g$ be the positive literal contained in $C$. For a suitable $(k-1)$-set of variables $A$, we have $C=\mathrm{neg}(A)\cup\{g,l\}$, $g$ and $l$ have the opposite sign, and $\alpha$ falsifies all the literals in $\mathrm{neg}(A)$ and $g$. Assume that the variable in $g$ is $x_{i}$ and the variable in $l$ is $x_{j}$. Hence, we have either $g=x_{i}$ and $l=\neg x_{j}$ or $g=\neg x_{i}$ and $l=x_{j}$. By assumption, $G^{d}(\varphi,A)$ contains paths between $x_{i}$ and $x_{j}$ in both directions. If $g=x_{i}$, then falsifiying $g$ means setting $x_{i}$ to 00 and Lemma 5.2 with $b=0$ implies that UCP derives $x_{j}=0$ and, hence, the literal $l$. If $g=\neg x_{i}$, then falsifiying $g$ means setting $x_{i}$ to $1$ and Lemma 5.2 with $b=1$ implies that UCP derives $x_{j}=1$ and, hence, the literal $l$.
We obtained that every clause $C\in\Psi_{n,k}\setminus\varphi$ is absorbed by $\varphi$, hence, $\Psi_{n,k}$ is absorbed by $\varphi$. Since $\varphi$ is a subset of $\Psi_{n,k}$, $\varphi$ is ucp-equivalent to $\Psi_{n,k}$. ∎
Proposition 5.4.
The size (number of clauses) of an arbitrary formula ucp-equivalent to $\Psi_{n,k}$ is at least $n\,C(n-1,k,k-1)$.
Proof.
Let $\varphi^{*}$ be a smallest formula ucp-equivalent to $\Psi_{n,k}$. If $C\in\varphi^{*}$ is not a prime implicate of $\varphi^{*}$, then replacing it by a prime subimplicate can only make UCP stronger. Since $\varphi^{*}$ is propagation complete, this replacement leads to a formula ucp-equivalent to $\varphi^{*}$ and $\Psi_{n,k}$. It follows that without loss of generality, we can assume that $\varphi^{*}$ consists only of prime implicates and, hence, is a subset of $\Psi_{n,k}$. For every variable $x\in X$, let
$\mathcal{S}_{x}=\{B\mid\mathrm{neg}(B)\vee x\in\varphi^{*}\}\;.$
Clearly, $\mathcal{S}_{x}$ is a system of $k$-sets $B\subseteq X\setminus\{x\}$.
Let $x\in X$ and let $A$ be an arbitrary $(k-1)$-subset of $X\setminus\{x\}$. By Proposition 5.3, vertex $x$ has non-zero in-degree in $G^{d}(\varphi^{*},A)$ and, hence, there is a set $B\in\mathcal{S}_{x}$ such that $A\subseteq B$. It follows that $\mathcal{S}_{x}$ is a system of $k$-subsets of an $(n-1)$-set covering all $(k-1)$-subsets. Hence, $|\mathcal{S}_{x}|\geq C(n-1,k,k-1)$ and, since $|\varphi^{*}|=\sum_{x\in X}|\mathcal{S}_{x}|$, the statement follows. ∎
For analysis of the hypergraph formulas $\theta(H)$ we use the following notion.
Definition 5.5.
We say that a $(k+1)$-uniform hypergraph $H$ on the set of vertices $X$ has connected restrictions, if for every $(k-1)$-set $A\subseteq X$, the graph
$G(H,A)=\{e\setminus A\mid A\subseteq e\in H\}$ (13)
is a connected graph on the set $X\setminus A$.
Proposition 5.6.
If $H$ is $(k+1)$-uniform hypergraph on $X$, then $\theta(H)$ is ucp-equivalent to $\Psi_{n,k}$ if and only if $H$ has connected restrictions.
Proof.
Since $\theta(H)$ is a subset of $\Psi_{n,k}$, we can use the characterization of formulas ucp-equivalent to $\Psi_{n,k}$ from Proposition 5.3.
Let $A$ be an arbitrary $(k-1)$-subset of $X$. If $G^{d}(\theta(H),A)$ contains the directed edge $(x_{i},x_{j})$, then it also contains the directed edge $(x_{j},x_{i})$. Moreover, these two edges are contained in $G^{d}(\theta(H),A)$ if and only if $G(H,A)$ contains the undirected edge $\{x_{i},x_{j}\}$. Hence, $G^{d}(\theta(H),A)$ is strongly connected if and only if $G(H,A)$ is connected. Since this is satisfied for arbitrary $(k-1)$-subset $A\subseteq X$, $\theta(H)$ satisfies the condition from Proposition 5.3 if and only if $H$ has connected restrictions. ∎
Note that $\Psi_{n,k}=\theta(H_{0})$, where $H_{0}$ is the complete $(k+1)$-uniform hypergraph. The ucp-irredundant formula $\varphi^{\ell}$ can be obtained as follows.
Definition 5.7.
Let $H_{1}$ consist of all $(k+1)$-subsets of $X$ containing the variable $x_{1}$ and let $\varphi^{\ell}=\theta(H_{1})$.
In particular, $H_{1}$ has connected restrictions, since for every $(k-1)$-set $A$ not containing $x_{1}$, the graph $G(H_{1},A)$ is a star on $X\setminus A$ whose central vertex is $x_{1}$. If $x_{1}\in A$, then $G(H_{1},A)$ is a complete graph on $X\setminus A$.
Proposition 5.8.
Formula $\varphi^{\ell}$ is ucp-irredundant and ucp-equivalent to $\Psi_{n,k}$.
Proof.
As mentioned above, $H_{1}$ has connected restrictions. It follows that $\varphi^{\ell}$ is ucp-equivalent to $\Psi_{n,k}$ by Proposition 5.6. Let us prove that it is ucp-irredundant. For this purpose, let $C\in\varphi^{\ell}$ be arbitrary and let us prove that $\varphi^{\ell}\setminus\{C\}$ is not ucp-equivalent to $\Psi_{n,k}$. Let $V$ be the set of variables used in $C$ except of $x_{1}$, so we have $|V|=k$. Let $\alpha$ be the partial assignment of the variables in $V$ such that all literals on variables from $V$ in $C$ are falsified. The clause $C$ together with $\alpha$ derives a value of $x_{1}$ which makes the literal on $x_{1}$ in $C$ satisfied. Since $C$ is an implicate of $\Psi_{n,k}$, every formula ucp-equivalent to $\Psi_{n,k}$ should derive the same value of $x_{1}$ as $C$. The proof will be finished by proving that $\varphi^{\ell}\setminus\{C\}$ does not satisfy this.
Depending on whether $C$ contains the literal $x_{1}$ or $\neg x_{1}$, the partial assignment $\alpha$ assigns all variables in $V$ the value $1$ or all variables in $V$ are assigned the value $1$ except of one which is assigned 00. In both cases, the following is true. Every clause $D$ in $\varphi^{\ell}\setminus\{C\}$ on the same set of variables as $C$ is satisfied by $\alpha$. The reason is that two of the variables in $C$ have different sign in $C$ and $D$, at least one of them is not $x_{1}$, and the literal with this variable is satisfied in $D$. Every clause in $\varphi^{\ell}\setminus\{C\}$ on a set of variables different from the set of variables of $C$ contains at least two variables not assigned by $\alpha$, namely $x_{1}$ and some variable not in $V\cup\{x_{1}\}$ and, hence, cannot be used for UCP. It follows that UCP on $\varphi^{\ell}\setminus\{C\}$ with the assignment $\alpha$ does not derive a value of any new variable, in particular, a value of $x_{1}$. ∎
Construction of a hypergraph $H$ such that $\varphi^{*}=\theta(H)$ is a small formula ucp-equivalent to $\Psi_{n,k}$ is more complicated and we use results proven in the next two sections for this purpose.
6 Random permutations of a graph or hypergraph
The purpose of this section is to prove that a union of independent permutations of several graphs without isolated vertices is connected with high probability. This is an auxiliary result used in the next section to prove that union of several independent random permutations of a $(k+1)$-uniform hypergraph which covers all $k$-subsets is a hypergraph with connected restrictions with high probability.
Definition 6.1.
Let $H$ be a hypergraph on a set of vertices $V$. Consider a permutation $\pi:V\to V$. We denote by $\pi(H)$ the hypergraph on the set of vertices $V$ such that for every set $e\subseteq V$, $\pi(e)$ is a hyperedge of $\pi(H)$ is and only if $e$ is a hyperedge of $H$.
Intuitively, hypergraph $\pi(H)$ is obtained by moving each vertex $u$ of $H$ to $\pi(u)$ and moving the edges correspondingly. By a random permutation of $H$, we mean $\pi(H)$, where $\pi$ is chosen at random from the uniform distribution on permutations of $V$.
The following consequence of LYM inequality [19] is an anticoncentration inequality for the sum of independent random variables $X_{i}$ distributed uniformly on $\{0,a_{i}\}$ where $a_{i}\geq 2$ are integers. For technical reasons, it is formulated using counts instead of probabilities.
Lemma 6.2 ([20]).
Assume integers $a_{i}\geq 2$ for $i=1,\ldots,r$ and let $m=\sum_{i=1}^{r}a_{i}$. Then for every $2\leq d\leq\lfloor m/2\rfloor$, the number of sets $I\subseteq\{1,\ldots,r\}$, such that
$\sum_{i\in I}a_{i}=d$ (14)
is at most
(${\lfloor m/2\rfloor\choose\lfloor d/2\rfloor}\;.$
Proof.
The sets $I$ satisfying (14) are pairwise incomparable and have size at most $\lfloor d/2\rfloor$. Let $c_{j}$ for $j=1,\ldots,\lfloor d/2\rfloor$ be the number of these sets of size $j$. By LYM inequality [19], we have
$\sum_{j=1}^{\lfloor d/2\rfloor}\frac{c_{j}}{{r\choose j}}\leq 1\;.$
Since $r\leq\lfloor m/2\rfloor$, we have also
$\sum_{j=1}^{\lfloor d/2\rfloor}\frac{c_{j}}{{\lfloor m/2\rfloor\choose j}}\leq 1$
and, since $\lfloor d/2\rfloor\leq\lfloor m/2\rfloor/2$, this implies that the number of sets $I$ satisfying (14) is at most
$\sum_{j=1}^{\lfloor d/2\rfloor}c_{j}\leq{\lfloor m/2\rfloor\choose\lfloor d/2\rfloor}$
as required. ∎
Let $G=(V,E)$ be an undirected graph. For a permutation $\pi:V\to V$, we define $\pi(G)$ in the same way as for a hypergraph. Namely, for every pair of distinct vertices $u,v\in V$, $\{\pi(u),\pi(v)\}$ is an edge of $\pi(G)$ if and only if $\{u,v\}$ is an edge of $G$.
Lemma 6.3.
Assume a set $Y$ of size $m$ and let $G$ be a graph on $Y$ with no isolated vertices. Let $A\subseteq Y$ such that $22\leq|A|\leq m/$ 2 and denote $d=|A|$. The probability that a random permutation of $G$ has no edge connecting $A$ and $Y\setminus A$ is at most
(${\lfloor m/2\rfloor\choose\lfloor d/2\rfloor}{m\choose d}^{-1}$
Proof.
Consider a random permutation $\pi$ of the vertices of $G$. Graph $\pi(G)$ has no edge between $A$ and $Y\setminus A$ if and only if the preimage $\pi^{-1}(A)$ is a union of connected components of $G$. The preimages $\pi^{-1}(A)$ are uniformly distributed over subsets of $Y$ of size $d$. The number of subsets of $Y$ of size $d$ that are a union of connected components in $G$ is equal to the number of solutions of the equation (14) where $r$ is the number of connected components of $G$ and $a_{i}$ are their sizes. Since $G$ has no isolated vertices, we have $a_{i}\geq 2$ for $i=1,\ldots,r$ and, hence, the lemma follows from Lemma 6.2. ∎
Proposition 6.4.
Let $G_{1},\ldots,G_{s}$, $s\geq 3$, be graphs on the same set of $m$ vertices each of which has no isolated vertices. If $s$ is fixed and $m$ tends to infinity, then the probability that the union of independent random permutations of $G_{1},\ldots,G_{s}$ is not a connected graph is at most $O(m^{2-s})$.
Proof.
Let $Y$ be the common set of vertices of the graphs. The probability of the event that the union of independent permutations of $G_{1},\ldots,G_{s}$ is not a connected graph is at most the sum of the probabilities over subsets $A\subseteq Y$ of size between $2$ and $m/2$ of the event that the union contains no edge between $A$ and $Y\setminus A$. Using $d=|A|$ as a summation index and using Lemma 6.3, this is at most
$\sum_{d=2}^{\lfloor m/2\rfloor}p_{d}$
where
$p_{d}={\lfloor m/2\rfloor\choose\lfloor d/2\rfloor}^{s}{m\choose d}^{1-s}\;.$ (15)
For a few smallest values of $d$, the following simple estimate is sufficient. If $d$ and $s$ are fixed and $m\to\infty$, we have
$p_{d}=O\left(m^{sd/2+(1-s)d}\right)=O\left(m^{(2-s)d/2}\right)\;.$ (16)
For larger values of $d$, we estimate $p_{d}$ using a more precise bound on binomial coefficients. Denote
$b(t,d)=\left(\frac{t}{d}\right)^{d}\left(\frac{t}{t-d}\right)^{t-d}=\mathrm{e}^{t\mathrm{H}(d/t)}$
where $\mathrm{H}(x)=-x\ln x-(1-x)\ln(1-x)$ is the entropy function. Since $\mathrm{H}(x)$ is increasing for $20\leq x\leq 1/$ 2, $b(t,d)$ is increasing in $d$ if $t$ is fixed and $21\leq d\leq t/$ 2. Since $\mathrm{H}(x)$ is strictly concave, $\mathrm{H}(x)/x$ is decreasing for $0<x\leq 1$ and, hence, $b(t,d)$ is increasing in $t$ if $d$ is fixed. By a well-known upper bound on binomial coefficients and using $d\leq\lfloor m/2\rfloor$, we get
(${\lfloor m/2\rfloor\choose\lfloor d/2\rfloor}\leq b(\lfloor m/2\rfloor,\lfloor d/2\rfloor)\leq b(m/2,d/2)=b(m,d)^{1/2}\;.$
By a well-known lower bound on binomial coefficients, we have
$\frac{1}{m+1}b(m,d)\leq{m\choose d}\;.$
Using these bounds and (15), we obtain
$p_{d}\leq(m+1)^{s-1}b(m,d)^{1-s/2}\;.$ (17)
Let $d_{0}=\lceil 4(s-1)/(s-2)\rceil$. Using (16) for $d\geq 2$, we obtain
$\sum_{d=2}^{d_{0}-1}p_{d}=O(m^{2-s})\;.$ (18)
If $d_{0}\leq d\leq m/2$, then $b(m,d)\geq b(m,d_{0})\geq(m/d_{0})^{d_{0}}$ and by (17)
$p_{d}\leq(m+1)^{s-1}(m/d_{0})^{d_{0}(1-s/2)}=O(m^{1-s})\;.$
Together with (18), we obtain
$\sum_{d=2}^{\lfloor m/2\rfloor}p_{d}=O(m^{2-s})$ (19)
and the proof is finished. ∎
7 Size of a hypergraph with connected restrictions
In this section, we prove an upper bound on the minimum size of a hypergraph with connected restrictions using covering numbers.
Proposition 7.1.
For every large enough $n$ and $k=n/2+O(1)$, there is a $(k+1)$-uniform hypergraph on $n$ vertices of size at most $6\,C(n,k+1,k)$ with connected restrictions.
Proof.
Let $|X|=n$ and let $H$ be a $(k+1)$-uniform hypergraph of size $C(n,k+1,k)$ covering all $k$-subsets of $X$. Let $s=5$, let $\pi_{1},\ldots,\pi_{s}$ be independent random permutations of $X$ chosen from the uniform distribution, and let $H^{*}=\pi_{1}(H)\cup\ldots\cup\pi_{s}(H)$.
If $A$ is a $(k-1)$-subset of $X$, then each of the graphs $G(\pi_{i}(H),A)$ has no isolated vertices and $G(H^{*},A)$ is their union, however, we cannot directly use Proposition 6.4, since $G(\pi_{i}(H),A)$ is not a random permutation of a fixed graph on $X\setminus A$. In order to overcome this, we partition the probability space of all $s$-tuples of permutations of $X$ by a set of mutually exclusive conditions such that the assumption of Proposition 6.4 is satisfied by the conditional distribution of the $s$-tuple of graphs $G(\pi_{i}(H),A)$, $i=1,\ldots,s$ under any of these conditions. Namely, for an arbitrary $s$-tuple $(A_{1}^{\prime},\ldots,A_{s}^{\prime})$ of $(k-1)$-subsets of $X$ let $c(A_{1}^{\prime},\ldots,A_{s}^{\prime})$ be the condition $\pi_{i}(A_{i}^{\prime})=A$ for all $i=1,\ldots,s$. Clearly, these conditions represent a partition of the probability space and all of them have the same probability.
Fix an arbitrary $s$-tuple $(A_{1}^{\prime},\ldots,A_{s}^{\prime})$ and consider the conditional distribution under $c(A_{1}^{\prime},\ldots,A_{s}^{\prime})$. For a fixed $i$ and all permutations $\pi_{i}$ satisfying $\pi_{i}(A_{i}^{\prime})=A$, the graphs $G(\pi_{i}(H),A)$ are isomorphic, since all of them are isomorphic to $G(H,A_{i}^{\prime})$. Choose any of the graphs $G(\pi_{i}(H),A)$ as $G_{i}$. The distribution of $G(\pi_{i}(H),A)$ is the same as the distribution of random permutations of $G_{i}$. Moreover, these random permutations are independent for $i=1,\ldots,s$. Hence, we can use Proposition 6.4 with $Y=X\setminus A$ to obtain that $G(H^{*},A)$ is not connected with probability at most $O((n-k+1)^{2-s})$ under any of the considered conditions and, hence, also unconditionally.
A $(k-1)$-set $A$ is called “bad”, if the graph $G(H^{*},A)$ is not connected and let $t$ be the random variable equal to the number of such sets. By the argument above, we have
$\mathbf{E}\;t=O\left((n-k+1)^{2-s}\right){n\choose k-1}\;.$ (20)
For each of the $t$ “bad” sets $A$, we add several $(k+1)$-sets to $H^{*}$ to make $G(H^{*},A)$ connected. The number of $(k+1)$-sets needed for this is at most the number of the connected components $G(H^{*},A)$ minus $1$ which is at most $(n-k-1)/2$. Hence, the resulting $(k+1)$-uniform hypergraph $H^{**}$ has size at most $s|H|+t(n-k-1)/2$. By (20), there is a choice of permutations $\pi_{i}$ for $i=1,\ldots,s$, such that
$t\leq O\left((n-k+1)^{2-s}\right){n\choose k-1}\leq O\left((n-k+1)^{2-s}\right){n\choose k}$
Since $n-k=\Theta(k)$, we can conclude
$|H^{**}|\leq s|H|+O\left(k^{3-s}\right){n\choose k}\;.$
Using (1) and $s=5$, we get
$|H^{**}|\leq(s+1)|H|\leq 6\,C(n,k+1,k)$
if $n$ and, hence, also $k$ are sufficiently large. This finishes the proof, since $H^{**}$ has connected restrictions. ∎
8 Proof of the results
In this section, we summarize the results of the previous sections. First, we finish the proof of Theorem 4.4 on a lower and an upper bound on the size of a smallest formula ucp-equivalent to $\Psi_{n,k}$.
Proof of Theorem 4.4.
The lower bound is Proposition 5.4. For the upper bound, consider a hypergraph $H$ of size at most $6\,C(n,k+1,k)$ with connected restrictions guaranteed by Proposition 7.1. By Proposition 5.6, $\theta(H)$ is ucp-equivalent to $\Psi_{n,k}$ and its size is $(k+1)|H|$. It follows that the size of a smallest formula ucp-equivalent to $\Psi_{n,k}$ is at most $6\,(k+1)C(n,k+1,k)$. Using (2) and the bound (5) or (7), we obtain the required asymptotic estimate. ∎
The main result is Theorem 4.1 providing an almost linear separation between the size of a specific ucp-irredundant formula ucp-equivalent to $\Psi_{n,k}$ and a smallest such formula.
Proof of Theorem 4.1.
Let $\varphi^{\ell}$ be the formula from Definition 5.7. This formula is an irredundant PC formula equivalent to $\Psi_{n,k}$ by Proposition 5.8. Assuming $k=n/2+O(1)$, its size is
$|\varphi^{\ell}|=(k+1){n-1\choose k}=(k+1)\left(1-\frac{k}{n}\right){n\choose k}=\Theta(n){n\choose k}\;.$
Let $\varphi^{*}$ be a smallest formula ucp-equivalent to $\Psi_{n,k}$. The statement follows by comparing the lower bound on $|\varphi^{\ell}|$ and the upper bound on $|\varphi^{*}|$ implied by Theorem 4.4. ∎
Let us include an observation concerning the gap between the lower and the upper bound in Theorem 4.4. These two bounds differ by a factor $\Theta(\mu(n,k)/\mu(n-1,k-1))$ where $\mu$ is defined in (2). The following proposition demonstrates that the density of the set of pairs $(n,k)$ satisfying $n=2k+O(1)$ and for which the ratio $\mu(n,k)/\mu(n-1,k-1)$ is bounded by a constant cannot decrease faster with $k$ than one over a doubly logarithmic function.
Proposition 8.1.
Let $a\in\mathbb{Z}$. For every large enough $k_{0}$, there is an integer $k$ satisfying $k_{0}+1\leq k\leq k_{0}+O(\ln\ln k_{0})$ and such that for $n=2k+a$ we have
$\mu(n-1,k-1)\leq\mu(n,k)\leq 3\,\mu(n-1,k-1)\;.$
Proof.
By (8), we have $\mu(n-1,k-1)\leq\mu(n,k)$ for every $n>k>1$. Assume $k_{0}\geq\max(1,5\,|a+1|)$, $t\geq 1$ and assume that for all $k_{0}+1\leq k\leq k_{0}+t$ and $n=2k+a$, we have
$\frac{\mu(n,k)}{\mu(n-1,k-1)}>3$ (21)
By (9) and for a suitable $c$, we have
$\frac{\mu(n+1,k)}{\mu(n,k)}\geq\frac{1}{2}\left(1+\frac{a+1}{2k+a+1}\right)\geq\frac{1}{2}\left(1-\frac{1}{9}\right)=c>\frac{1}{3}$
and together with (21), we obtain for every $k_{0}+1\leq k\leq k_{0}+t$ and $n=2k+a$
$1<3c<\frac{\mu(n+1,k)}{\mu(n-1,k-1)}=\frac{\mu(2k+1+a,k)}{\mu(2k-1+a,k-1)}$
Multiplying this for $k_{0}+1\leq k\leq k_{0}+t$, simplifying the telescoping product, and using (5), we obtain
$(3c)^{t}<\frac{\mu(2k_{0}+2t+1+a,k_{0}+t)}{\mu(2k_{0}+1+a,k_{0})}\leq 1+\ln(k_{0}+t+1)$
Choosing, for example,
$t=\frac{\ln\ln(4k_{0})}{\ln(3c)}$
we get a contradiction if $k_{0}$ is sufficiently large and this implies the statement. ∎
9 Conclusion and further research
The paper introduces ucp-equivalence on CNF formulas and investigates the ratio of the size of a ucp-irredundant formula and the size of a smallest ucp-equivalent formula. The paper shows that a general upper bound on this ratio is at most $n^{2}$ (Proposition 3.6) and at least $\Omega(n/\ln n)$ (Theorem 4.1). The gap between these bounds is quite large. An interesting open problem for further research is to reduce this gap.
Proposition 8.1 demonstrates that the lower and the upper bound on the size of a smallest formula ucp-equivalent to $\Psi_{n,k}$ in Theorem 4.4 differ at most by a multiplicative constant for a set of pairs $(n,k)$ satisfying $n=2k+O(1)$ and whose density tends to zero with $k$ not faster than one over a doubly logarithmic function. An interesting question is whether the lower and the upper bound in Theorem 4.4 differ at most by a multiplicative constant for all pairs $(n,k)$ satisfying $n=2k+O(1)$.
Acknowledgement. The author would like to thank Diana Piguet for a fruitful discussion on using hypergraphs, in particular, for suggesting Definition 4.5, and Matas Šileikis for suggesting Lemma 6.2 which improves a slightly weaker bound on the same quantity by the author. The author acknowledges institutional support RVO:67985807.