### v5

# On CNF formulas irredundant with respect to unit clause propagation

Abstract
Abstract
Two CNF formulas are called ucp-equivalent, if they behave in the same way with respect to the unit clause propagation (UCP). A formula is called ucp-irredundant, if removing any clause leads to a formula which is not ucp-equivalent to the original one. As a consequence of known results, the ratio of the size of a ucp-irredundant formula and the size of a smallest ucp-equivalent formula is at most $n^{2}$, where $n$ is the number of the variables. We demonstrate an example of a ucp-irredundant formula for a symmetric definite Horn function which is larger than a smallest ucp-equivalent formula by a factor $\Omega(n/\ln n)$. Consequently, a general upper bound on the above ratio cannot be smaller than this.
1 Introduction
In some contexts, the strength of unit clause propagation (UCP) on the formula is important and formulas representing the same function have to be distinguished according to the behavior of UCP on them. An important example is the set of learned clauses in a SAT solver. Most of these clauses are implicates of the original formula so they do not change the represented function and they are added in order to make UCP stronger. In order to investigate CNF formulas in such a context, we need to compare the strength of UCP on them. In this paper, we consider a corresponding equivalence and irredundancy.
Two CNF formulas are called ucp-equivalent, if for every partial assignment $\alpha$, UCP derives the same consistent set of literals from the two formulas together with $\alpha$ or we get a contradiction in both cases. Considering total assignments $\alpha$, one can verify that ucp-equivalent formulas represent the same function, so they are equivalent.
Let us present a few examples demonstrating the notion of ucp-equivalence. Formulas
$\varphi_{1}$ $=$ $(a\vee c)\wedge(b\vee c)$
$\varphi_{2}$ $=$ $(a\vee c)\wedge(\neg a\vee b\vee c)$
are equivalent, but not ucp-equivalent, since UCP derives the literal $c$ from the formula $\varphi_{1}\wedge\neg b$ and not from the formula $\varphi_{2}\wedge\neg b$. Formula $\varphi_{2}$ is not prime, since the literal $\neg a$ can be removed without changing the represented function. Consider prime formulas
$\varphi_{3}$ $=$ ($(\neg a\vee b)\wedge(\neg a\vee c)\wedge(\neg b\vee\neg c\vee d)$
$\varphi_{4}$ $=$ $\varphi_{3}\wedge(\neg a\vee d)$
which are equivalent, since $\neg a\vee d$ is a consequence of $\varphi_{3}$. These formulas are not ucp-equivalent, since UCP on $\varphi_{4}\wedge\neg d$ derives $\neg a$ while not on $\varphi_{3}\wedge\neg d$.
Ucp-equivalent formulas have to be much closer to each other than equivalent formulas. However, ucp-equivalent formulas may not contain any common clause. An example of a pair of such formulas is
$\varphi_{5}$ $=$ ($(\neg a\vee b)\wedge(\neg b\vee c)\wedge(\neg c\vee a)$
$\varphi_{6}$ $=$ ($(\neg a\vee c)\wedge(\neg c\vee b)\wedge(\neg b\vee a)\;.$
Without an additional partial assignment, UCP derives no new literal on both these formulas. After adding a positive literal on any of the variables $a$, $b$, or $c$, UCP derives all the remaining positive literals on both the formulas and the same is true for the negative literals. More generally, consider any finite set of variables and two conjunctions of implications between pairs of these variables which both form a strongly connected directed graph. An implication $x\to y$ is equivalent to the clause $\neg x\vee y$, so the conjunctions can be considered as CNF formulas. Using the same argument as above, such formulas are always ucp-equivalent, although the graphs can be quite different.
A formula is called ucp-irredundant, if removing any clause changes the represented function or makes UCP on the formula weaker. In both these cases, the obtained formula is not ucp-equivalent to the original one. All the formulas $\varphi_{1}$, …, $\varphi_{6}$ are ucp-irredundant. Removing any clause from them except of $\neg a\vee d$ from $\varphi_{4}$ changes the represented function. Removing the clause $\neg a\vee d$ does not change the function represented by $\varphi_{4}$, but makes UCP on the formula weaker.
UCP on a CNF formula $\varphi$ of $n$ variables can be precisely represented using a specific Horn formula on $2n$ variables which we call implicational dual rail encoding of $\varphi$ and denote $IDR(\varphi)$. The construction is a Horn part of a formula used in [1, 2] to simulate UCP. The variables of this formula represent literals on the variables of the original formula. Such a representation is known as a dual rail encoding, see, for example [3, 4, 5], where it is used to reduce SAT problem to MaxSAT for a specific Horn formula. The implicational dual rail encoding of a formula $\varphi$ is also a Horn formula, however, it is quite different. It contains $k$ definite Horn clauses for every clause of length $k$ in $\varphi$, since each of the $k$ possible steps of UCP in the clause is represented by a separate Horn clause or, equivalently, by an implication.
Let us point out that representation of UCP using a Horn formula is used in [1] to characterize the size of a CNF decomposition of a propagator known also as a CNF encoding satisfying generalized arc consistency for an arbitrary constraint using the size of a specific monotone Boolean circuit up to a polynomial relation. In [2] a formula whose Horn part represents UCP in some other formula is used to construct a CNF encoding of a disjunction of several CNF formulas preserving the strength of UCP of the input formulas. In [6], implicational dual rail encoding is used to construct an algorithm computing a propagation complete formula using results on learning definite Horn formulas.
Two formulas are ucp-equivalent if and only if their implicational dual rail encodings are equivalent in the sense to represent the same Horn function. It is a well-known result [7] that the ratio of the size of a Horn formula irredundant with respect to representing a given function and the size of a smallest equivalent formula is at most the number of the variables.
Situations when the size of an irredundant and a smallest equivalent structure are close to each other are rare in complexity. In particular, this is not true for CNF formulas with respect to equivalence. An unsatisfiable formula of $n$ variables consisting of all $2^{n}$ clauses of length $n$ is minimally unsatisfiable, since removing any of its clauses makes it satisfiable. Hence, this formula is irredundant in the sense that removing any clause changes the represented function. However, the smallest equivalent formula consists only of an empty clause.
Horn formulas represent a very narrow class of Boolean functions. Using the implicational dual rail encoding we can transform the result for Horn formulas to ucp-irredundant formulas representing an arbitrary Boolean function. In this case, we consider the ratio of the size of any ucp-irredundant formula and the size of a smallest ucp-equivalent formula. Since implicational dual rail encoding changes the number of the clauses, we get a bound which is worse than in the case of Horn formulas. Namely, the upper bound is $2n^{2}$, where $n$ is the number of the variables. A slightly better bound $n^{2}$ can be obtained by a proof used in [8] to prove the same bound for propagation complete formulas since the assumption of propagation completeness is used only to guarantee ucp-equivalence. Moreover, it appears that the factor $n^{2}$ holds both for the number of clauses and the length of the formula.
An interesting question is what is the maximum value of the ratio of the size of any ucp-irredundant formula and the size of a smallest ucp-equivalent formula in terms of the number of the variables. It is plausible to assume that the upper bound $n^{2}$ can be improved, however, this is an open problem. The main result of the paper is an example of a ucp-irredundant formula $\varphi^{\ell}$ of $n$ variables which is larger than the smallest ucp-equivalent formula $\varphi^{*}$ by a factor $\Omega(n/\ln n)$. It follows that a general upper bound on the ratio cannot be smaller than this.
The formulas used in the example represent a symmetric definite Horn function and the construction of the formula $\varphi^{\ell}$ is easy. The hard part of the proof is to derive an upper bound on the size of $\varphi^{*}$ and we show that the size of $\varphi^{*}$ is closely related to the size of a specific covering system. Namely, for $n$ and $k=n/2+O(1)$, we use a system of $(k+1)$-subsets of an $n$-set such that every $k$-set is contained in at least one member of the system. The proof of the needed properties of such systems and their relationship to the formulas uses probabilistic method and takes a significant part of the paper. The proof is finished using a known upper bound on the size of such a system from [9] or a different upper bound from [10, 11]. It is an interesting question whether the lower bound $\Omega(n/\ln n)$ or an even better bound can be proven using a different type of functions and without the use of probabilistic method.
Section 3 defines ucp-equivalence and ucp-irredundancy and Section 4 formulates the results. The property of the formulas needed for the results is reformulated in Section 5 as a property of a hypergraph. The construction of the required hypergraph is described in Section 7 using an auxiliary result from Section 6. The proofs of the results are summarized in Section 8. Conclusion and questions for further research are formulated in Section 9.
2 Preliminaries
2.1 CNF formulas
A literal is a Boolean variable or its negation, a clause is a disjunction of literals on different variables, and a CNF formula is a conjunction of clauses. An empty clause denoted $\bot$ is a clause containing no literals and represents a contradiction. A unit clause is a clause containing exactly one literal. Since we consider only CNF formulas, we call them formulas for simplicity.
A formula is considered as a set of clauses and a clause is considered as a set of literals. In particular, we use set operations on formulas and clauses. If $A$ is a set of variables, let $\mathrm{neg}(A)=\{\neg x\mid x\in A\}$ be a set of literals used typically as a part of a clause. The size $|\varphi|$ of a formula $\varphi$ is the number of its clauses. The length of a clause is the number of the literals in it. The length $\|\varphi\|$ of a formula $\varphi$ is the sum of the lengths of its clauses.
A clause $C$ is an implicate of a formula $\varphi$, if every assignment of the variables satisfying $\varphi$ satisfies $C$. A clause $C$ is a prime implicate of a formula $\varphi$, if it is an implicate of $\varphi$ and no proper subset of $C$ is an implicate of $\varphi$.
A formula is a Horn formula, if every clause contains at most one positive literal, and a formula is a definite Horn formula, if every clause contains exactly one positive literal.
A partial assignment of the variables is a partial map from the variables to the set $\{0,1\}$ understood as a conjunction of identities of the form $x=a$, where $x$ is a variable and $a\in\{0,1\}$. A partial assignment will be represented by a conjunction of the literals equivalent to these identities. For example, the conjunction of identities $x_{1}=1\wedge x_{2}=0$ is equivalent to the conjunction of literals $x_{1}\wedge\neg x_{2}$.
Unit clause propagation (UCP) is a process of deriving literals or a contradiction implied by the formula using the following process. Assume a formula $\varphi$ and let us initialize a set $U$ as the set of clauses of length at most one contained in $\varphi$. Clauses of length one will be identified with the literals contained in them. All clauses in the initial set $U$ are elements of $\varphi$, so they are clearly consequences of $\varphi$. Assume a clause $C\in\varphi$, and a literal $l\in C$. If $U$ contains the negations of all the literals in $C\setminus\{l\}$, then every satisfying assignment of $\varphi$ satisfies $l$ and, hence, $l$ is a consequence of $\varphi$. Similarly, if $U$ contains the negation of all the literals in $C$, then $\bot$ is a consequence of $\varphi$. Unit clause propagation on a formula $\varphi$ is the process of iteratively extending $U$ by literals and possibly $\bot$ derived as consequences of this simple type. The process terminates if no further literal or $\bot$ can be derived or if $\bot\in U$.
Definition 2.1.
For every CNF formula $\varphi$, let $\mathrm{UCP}(\varphi)$ be
  • the set of literals contained in $\varphi$ as unit clauses or derived by unit clause propagation on $\varphi$, if the contradiction is not reached,
[…]
One can show that $\mathrm{UCP}(\varphi)$ does not depend on the order of the derivations performed by the process of UCP. If $\mathrm{UCP}(\varphi)=\{\bot\}$, then $\varphi$ is unsatisfiable. The opposite is not true. For example, the formula $\varphi_{0}=(x_{1}\vee x_{2})\wedge(x_{1}\vee\neg x_{2})\wedge(\neg x_{1}\vee x_{2})\wedge(\neg x_{1}\vee\neg x_{2})$ is unsatisfiable, however, $\mathrm{UCP}(\varphi_{0})=\emptyset$.
A non-empty partial assignment $\alpha$ can be used to simplify a formula $\varphi$ to a formula $\varphi^{\prime}$ as follows. For every $l\in\alpha$, every clause containing $l$ is removed and the negated literal $\neg l$ is removed from all other clauses. One can verify that $\mathrm{UCP}(\varphi\wedge\alpha)=\mathrm{UCP}(\varphi^{\prime}\wedge\alpha)$.
2.2 Covering and Turán numbers
A $k$-set or a $k$-subset is a set or a subset of size $k$. We use known bounds on covering numbers $C(n,r,k)$ [9, 12] and the closely related Turán numbers $T(n,\ell,t)$ [11, 10]. Since there are different conventions to define these numbers, we recall the definitions which we use.
Definition 2.2.
For integers $n>r>k>0$, let the covering number $C(n,r,k)$ be the smallest size of a system of $r$-subsets of an $n$-set such that every $k$-subset is contained in at least one set in the system.
Definition 2.3.
For integers $n>\ell>t>0$, let the Turán number $T(n,\ell,t)$ be the smallest size of a system of $t$-subsets of an $n$-set such that every $\ell$-subset contains at least one set in the system.
Taking complements of the sets, we obtain $C(n,r,k)=T(n,n-k,n-r)$. Note that Turán’s theorem implies a precise value of $T(n,\ell,2)$, if we minimize the number of edges to be removed from a complete graph on $n$ vertices in order to eliminate all complete subgraphs on $\ell$ vertices.
Let us recall several known results on covering numbers in the case $r=k+1$ which we use. Clearly,
$C(n,k+1,k)\geq\frac{1}{k+1}{n\choose k}\;.$ (1)
If $k$ is fixed and $n\to\infty$, then by [13], we have
$C(n,k+1,k)=(1+o(1))\frac{1}{k+1}{n\choose k}\;.$
We are mainly interested in cases when $k=n/2+O(1)$ for which only less precise bounds are known. Let $\mu(n,k)$ be such that
$C(n,k+1,k)=\frac{1}{k+1}{n\choose k}\mu(n,k)$ (2)
and, equivalently,
$T(n,t+1,t)=\frac{1}{t+1}{n\choose t}\mu(n,n-t-1)\;.$ (3)
For an upper bound on $\mu(n,k)$, let us recall the following.
Theorem 2.4 (a special case of Theorem 13.4 from [9]).
For every $n>k\geq 2$, we have
$C(n,k+1,k)\leq\frac{1}{k+1}{n\choose k}(1+\ln(k+1))$ (4)
Together with (1), this implies
$1\leq\mu(n,k)\leq 1+\ln(k+1)$ (5)
for every $n>k\geq 2$. An incomparable upper bound on $\mu(n,k)$ follows from the bound [10] on Turán numbers which improves [11] by a factor of $2$. By this bound, we have
$T(n,t+1,t)\leq\frac{(1+o(1))\ln t}{2t}{n\choose t}$ (6)
for $n\to\infty$ and $t\to\infty$. Combining (6) with (3), substituting $k=n-t-1$ and assuming $n-k\to\infty$, we obtain
$\mu(n,k)\leq\frac{1+o(1)}{2}\ln(n-k-1)\;.$ (7)
Bounds (5) and (7) are complementary in the sense that (5) is better, if $k$ is small, while (7) is better, if $k$ is close to $n$. Although the two bounds are incomparable in general, they are close to each other for $n$ and $k=n/2+O(1)$ used in this paper.
In Section 8 it is advantageous to rely on inequalities for the function $\mu(n,k)$ instead of inequalities for the covering numbers. The quantity $\mu(n,k)$ is logarithmically bounded and using it allows us to consider this relatively small factor independently of the main terms of the asymptotic estimate. Namely, we will need the following.
Lemma 2.5.
For every $n>k>1$, we have
$\mu(n-1,k-1)$$\leq$ $\mu(n,k)$ (8)
$\frac{n-k+1}{n+1}\mu(n,k)$$\leq$ $\mu(n+1,k)$ (9)
Proof.
Both inequalities are direct consequencies of inequalities for the covering numbers. In particular, by [14, 12], we have
$\frac{n}{k+1}C(n-1,k,k-1)\leq C(n,k+1,k)$
which implies (8) using (2).
Assume a system of $(k+1)$-subsets of a set $X$ of $n+1$ points of size $C(n+1,k+1,k)$ and covering all $k$-subsets. For any $x\in X$, one can create a system of $(k+1)$-subsets of $X\setminus\{x\}$ covering all $k$-subsets by replacing each occurence of $x$ in a $(k+1)$-tuple in the system by any element of $X\setminus\{x\}$ not contained in the $(k+1)$-tuple. Hence, we have
$C(n,k+1,k)\leq C(n+1,k+1,k)$
implying (9) using (2). ∎
3 Ucp-equivalence and ucp-irredundancy
Let us introduce the following equivalence relation on CNF formulas.
Definition 3.1.
Formulas $\varphi_{1}$ and $\varphi_{2}$ are ucp-equivalent if for every partial assignment $\alpha$, we have $\mathrm{UCP}(\varphi_{1}\wedge\alpha)=\mathrm{UCP}(\varphi_{2}\wedge\alpha)$.
By considering total assignments represented by $\alpha$ containing exactly one literal (positive or negative) on each variable, one can prove the following.
Proposition 3.2.
If two formulas are ucp-equivalent, then they are equivalent in the sense to represent the same function.
The notion of ucp-equivalence appeared implicitly in analysis of propagation complete (PC) formulas [15, 16], so let us recall their definition. A formula $\varphi$ is called propagation complete, if for every partial assignment $\alpha$ of the variables, UCP on $\varphi\wedge\alpha$ derives a contradiction, if $\varphi\wedge\alpha$ is inconsistent, and all literals implied by $\varphi\wedge\alpha$ otherwise. Since UCP is logically sound, a formula is propagation complete, if the strength of UCP on it is maximal among all formulas representing the same function. In particular, every two equivalent PC formulas are ucp-equivalent, since UCP on them is controlled by the function they represent. One can also verify that a formula is PC if and only if it is ucp-equivalent to the set of all its (prime) implicates.
Let us recall the notion of an absorbed clause introduced in [15]. Negation of a clause is the conjunction of the negations of its literals. A clause $C$ is absorbed by a formula $\varphi$, if for every $l\in C$, UCP on the formula $\varphi\wedge\neg(C\setminus\{l\})$ derives $l$ or a contradiction. A formula is called minimal PC formula in [15] if none of its clauses is absorbed by the remaining ones. This does not guarantee that no smaller equivalent PC formula exists, however, the formula is irredundant with respect to the behavior of UCP. We use a straightforward generalization of this type of irredundancy obtained by removing the requirement that the formula is propagation complete.
Definition 3.3.
A formula $\varphi$ is ucp-irredundant if any of the following equivalent conditions is satisfied:
  • None of the clauses $C\in\varphi$ is absorbed by $\varphi\setminus\{C\}$.
  • For every $C\in\varphi$, the formulas $\varphi$ and $\varphi\setminus\{C\}$ are not ucp-equivalent.
The notion of absorption can naturally be extended to formulas.
Definition 3.4.
A formula $\varphi_{1}$ is absorbed by the formula $\varphi_{2}$, which will be denoted $\varphi_{1}\leq_{ucp}\varphi_{2}$, if every clause of $\varphi_{1}$ is absorbed by $\varphi_{2}$.
One can verify that $\varphi_{1}\leq_{ucp}\varphi_{2}$ if and only if for every partial assignment $\alpha$, we have $\mathrm{UCP}(\varphi_{1}\wedge\alpha)\subseteq\mathrm{UCP}(\varphi_{2}\wedge\alpha)$ or $\mathrm{UCP}(\varphi_{2}\wedge\alpha)=\{\bot\}$. It follows that the relation $\leq_{ucp}$ is a preorder and any two formulas $\varphi_{1}$ and $\varphi_{2}$ are ucp-equivalent if and only if $\varphi_{1}\leq_{ucp}\varphi_{2}$ and $\varphi_{2}\leq_{ucp}\varphi_{1}$. In particular, $\leq_{ucp}$ defines a partial order on the classes of ucp-equivalent formulas and the class of PC formulas is the maximal class in this order. Since, by definition, $\varphi_{1}\leq_{ucp}\varphi_{2}$ can be tested in polynomial time, also ucp-equivalence can be tested in polynomial time.
By the result [8], if $\varphi$ is a ucp-irredundant PC formula, then the ratio of its size and the size of a smallest PC formula equivalent to $\varphi$ is at most $n^{2}$, where $n$ is the number of the variables. The assumption that the two formulas are PC is used only to guarantee that they behave in the same way with respect to UCP or, equivalently, that the two formulas are ucp-equivalent. It follows that the upper bound on the ratio of their sizes can be generalized to non-PC formulas as follows. The statement is the strongest if $\varphi^{*}$ is the smallest formula ucp-equivalent to $\varphi$, however, the proof does not use this assumption.
Proposition 3.5.
If $\varphi$ is a ucp-irredundant formula on $n$ variables and $\varphi^{*}$ is ucp-equivalent to $\varphi$, then
$|\varphi|\leq n^{2}|\varphi^{*}|$ (10)
and
$\|\varphi\|\leq n^{2}\|\varphi^{*}\|$ (11)
Proof.
Every clause of $\varphi^{*}$ is absorbed by $\varphi$. Let $\varphi^{\prime}$ be the set of clauses in $\varphi$ needed for this for all clauses of $\varphi^{*}$. For each pair $l\in C\in\varphi^{*}$, we need at most $n$ clauses of $\varphi$ to derive $l$ or a contradiction from $\varphi\wedge\neg(C\setminus\{l\})$. This implies $|\varphi^{\prime}|\leq n\|\varphi^{*}\|$. Since $\varphi^{\prime}\subseteq\varphi$, we have
$\varphi\leq_{ucp}\varphi^{*}\leq_{ucp}\varphi^{\prime}\leq_{ucp}\varphi$
which implies that the formula $\varphi^{\prime}$ is ucp-equivalent to $\varphi$. Since $\varphi$ is ucp-irredundant, $\varphi^{\prime}$ is equal to $\varphi$. Hence, we have $|\varphi|\leq n\|\varphi^{*}\|$ implying
$|\varphi|\leq n\|\varphi^{*}\|\leq n^{2}|\varphi^{*}|$
and
$\|\varphi\|\leq n|\varphi|\leq n^{2}\|\varphi^{*}\|$
as required. ∎
Using part of a construction used in [1, 2], UCP on a CNF formula can be represented by the following Horn formula. Assume, $\varphi=C_{1}\wedge\ldots\wedge C_{m}$ is a formula on variables $X=\{x_{1},\ldots,x_{n}\}$. For every literal $l$ on a variable from $X$ introduce a variable $\llbracket l\rrbracket$. A total assignment of these variables satisfying $\neg\llbracket x_{i}\rrbracket\vee\neg\llbracket\neg x_{i}\rrbracket$ for every variable $x_{i}\in X$ encodes the set of the satisfied literals under a partial assignment of the variables $X$. The implicational dual rail encoding of $\varphi$ is
$\mathrm{IDR}(\varphi)=\mathrm{IDR}(C_{1})\wedge\ldots\wedge\mathrm{IDR}(C_{m})\wedge\bigwedge_{i=1}^{n}(\neg\llbracket x_{i}\rrbracket\vee\neg\llbracket\neg x_{i}\rrbracket)$
where for a clause $C=l_{1}\vee\ldots\vee l_{k}$,
$\mathrm{IDR}(C)=\bigwedge_{i=1}^{k}\left(\bigwedge_{j\not=i}\llbracket\neg l_{j}\rrbracket\to\llbracket l_{i}\rrbracket\right)$
The set of models of $\mathrm{IDR}(\varphi)$ is precisely the set of encodings of partial assignments of the variables $X$ which are closed under UCP. This implies the following.
Proposition 3.6.
Two formulas $\varphi_{1}$ and $\varphi_{2}$ are ucp-equivalent if and only if the Horn formulas $\mathrm{IDR}(\varphi_{1})$ and $\mathrm{IDR}(\varphi_{2})$ are equivalent.
4 The results
The main result is the following separation between the size of a ucp-irredundant formula and a smallest ucp-equivalent formula.
Theorem 4.1.
For every sufficiently large $n$ and an integer $k=n/2+O(1)$, there is a ucp-irredundant formula $\varphi^{\ell}$ of $n$ variables and of size
$|\varphi^{\ell}|=(k+1){n-1\choose k}=\Theta(n){n\choose k}$
and such that a smallest ucp-equivalent formula $\varphi^{*}$ has size
$|\varphi^{*}|\leq 6(k+1)C(n,k+1,k)=O(\ln n){n\choose k}$
In particular, $|\varphi^{\ell}|/|\varphi^{*}|=\Omega(n/\ln n)$.
For a proof, we investigate the size of ucp-irredundant formulas ucp-equivalent to the following definite Horn formula $\Psi_{n,k}$.
Definition 4.2.
For any integers $n$, $k$, such that $11\leq k\leq n-$ 1, let $\Psi_{n,k}$ be the CNF formula consisting of all clauses on variables $X=\{x_{1},\ldots,x_{n}\}$ containing $k$ negative literals and one positive literal.
Clearly, $|\Psi_{n,k}|=(n-k){n\choose k}$. A Boolean function is called symmetric, if it is invariant under any permutation of its variables. Note that formula $\Psi_{n,k}$ represents the symmetric Boolean function $f_{n,k}$ of $n$ variables defined by
$f_{n,k}(x_{1},\ldots,x_{n})=1\Leftrightarrow\left(\sum_{i=1}^{n}x_{i}<k\right)\vee\left(\sum_{i=1}^{n}x_{i}=n\right)$ (12)
and, moreover, $\Psi_{n,k}$ is the set of all prime implicates of this function.
The proof of Theorem 4.1 relies on the upper bound on the size of a smallest formula ucp-equivalent to $\Psi_{n,k}$ implied by the following proposition. The lower bound on this size is included in order to demonstrate that the upper bound cannot be improved more than by a logarithmic factor.
Proposition 4.3.
For every $11\leq k\leq n-$ 1, the minimum size (number of clauses) of a CNF formula ucp-equivalent to $\Psi_{n,k}$ is at least
$n\,C(n-1,k,k-1)={n\choose k}\mu(n-1,k-1)\geq{n\choose k}$
and if $k=n/2+O(1)$ then it is at most
$6(k+1)C(n,k+1,k)=6\,{n\choose k}\mu(n,k)=O(\ln n){n\choose k}\;.$
The proof of Theorem 4.1 and of Proposition 4.3 are summarized in Section 8 using constructions presented in this and the following sections. The lower bound in Proposition 4.3 is proven in Section 5. For the proof of the remaining statements, we construct for infinitely many pairs $(n,k)$ ucp-irredundant formulas $\varphi^{\ell}$ (Section 5) and $\varphi^{*}$ (Section 8) ucp-equivalent to $\Psi_{n,k}$, such that $\varphi^{\ell}$ is as large as possible and $\varphi^{*}$ is as small as possible. The formula $\varphi^{*}$ which meets the upper bound in Proposition 4.3 and the formula $\varphi^{\ell}$ are then used to prove Theorem 4.1.
Each of the formulas $\varphi^{\ell}$ and $\varphi^{*}$ will be constructed as a hypergraph formula which simplifies their construction considerably. The lower bound in Proposition 4.3 is valid for an arbitrary subset of $\Psi_{n,k}$. This implies that using a hypergraph formula $\varphi^{*}$ cannot worsen the upper bound on $|\varphi^{*}|$ by more than a logarithmic factor compared to the minimum size for unrestricted formulas.
Definition 4.4 ([17, 18]).
For every $(k+1)$-uniform hypergraph $H$ whose vertices are the variables $X$, the hypergraph formula $\theta(H)$ consists of the clauses $\mathrm{neg}(e\setminus\{x\})\cup\{x\}$ for all hyperedges $e\in H$ and all variables $x\in e$.
Clearly, for every $H$ as above, $\theta(H)\subseteq\Psi_{n,k}$ and $|\theta(H)|=(k+1)|H|$.
5 A combinatorial characterization
In order to construct the formula $\varphi^{*}$, we are looking for a formula ucp-equivalent to $\Psi_{n,k}$ and as small as possible. According to Lemma 5.1 below, we can consider only prime formulas (containing only prime implicates) without loss of generality. Since the formula $\Psi_{n,k}$ is equal to the set of all its prime implicates, such formulas are subsets of $\Psi_{n,k}$. We present two combinatorial characterizations of the subsets of $\Psi_{n,k}$ which are ucp-equivalent to $\Psi_{n,k}$. The first of them is valid for an arbitrary subset of $\Psi_{n,k}$ and is used for a lower bound on the size of an arbitrary formula ucp-equivalent to $\Psi_{n,k}$. The other characterization is valid for subsets of $\Psi_{n,k}$ in the form of a hypergraph formula $\theta(H)$ defined above and is used later for a corresponding upper bound.
Lemma 5.1.
If $\varphi$ is an arbitrary CNF formula ucp-equivalent to $\Psi_{n,k}$, then there is a prime formula $\varphi^{\prime}$ ucp-equivalent to $\Psi_{n,k}$, such that $|\varphi^{\prime}|\leq|\varphi|$ and $\|\varphi^{\prime}\|\leq\|\varphi\|$.
Proof.
Let $\varphi$ be ucp-equivalent to $\Psi_{n,k}$. By Proposition 3.2, $\varphi$ represents the same function as $\Psi_{n,k}$, so these two formulas have the same set of prime implicates. Moreover, $\Psi_{n,k}$ is constructed as the set of all prime implicates of the function (12). Hence, $\Psi_{n,k}$ is propagation complete and, by assumption, also $\varphi$ is propagation complete. This means that UCP on $\varphi$ is the strongest possible among formulas representing the function (12).
If $C\in\varphi$ is not a prime implicate of $\varphi$, then replacing it by a prime subimplicate can only make UCP stronger. Since $\varphi$ is propagation complete, this replacement leads to a formula ucp-equivalent to $\varphi$ and $\Psi_{n,k}$. Repeating this, we replace all non-prime clauses of $\varphi$ by prime implicates and possibly remove redundant clauses. The resulting formula $\varphi^{\prime}$ satisfies the statement. ∎
Every clause of $\Psi_{n,k}$ contains $k$ negative literals. A partial assignment setting at most $k-2$ variables to $1$ and possibly some variables to 00 simplifies $\Psi_{n,k}$ to a formula where every clause contains at least $2$ negative literals. This implies that UCP on $\Psi_{n,k}$ and such a partial assignment does not derive any new literal and the same is true for every subset of $\Psi_{n,k}$. It follows that when considering UCP on subsets of $\Psi_{n,k}$ it is sufficient to consider partial assignments setting at least $k-1$ variables to $1$ and possibly some variables to 00.
Moreover, it appears that in order to characterize subsets $\varphi$ of $\Psi_{n,k}$ that are ucp-equivalent to $\Psi_{n,k}$, it is reasonable to split such an assignment into an initial setting of exactly $k-1$ variables to $1$, determine the set of binary clauses in the corresponding simplified formula, and then consider a further setting of one other variable, say $x_{\ell}$, to 00 or $1$. Depending on whether $x_{\ell}$ is set to 00 or $1$, UCP on $\varphi$ should derive all negative or all positive literals on the remaining variables, since this is true for UCP on the formula $\Psi_{n,k}$.
The binary clauses obtained by simplification after setting of exactly $k-1$ variables to $1$ necessarily contain one positive and one negative literal and, hence, represent an implication. We represent this set of implicatins by the following directed graph.
Definition 5.2.
Assume $\varphi\subseteq\Psi_{n,k}$ and a $(k-1)$-set $A\subseteq X$. Let $G^{d}(\varphi,A)$ be the directed graph on $X\setminus A$ defined as
$G^{d}(\varphi,A)=\{(x_{i},x_{j})\mid\mathrm{neg}(A)\cup\{\neg x_{i},x_{j}\}\in\varphi\}\;.$
Let $\varphi$ and $A$ be as in Definition 5.2 and let $\alpha=\bigwedge_{x\in A}x$ be a partial assignment setting all variables in $A$ to $1$. Then UCP on $\varphi\wedge\alpha$ derives no additional literals, since $\alpha$ extended by literals of the same polarity (positive or negative) on all the remaining variables is a satisfying assignment of $\Psi_{n,k}$ and $\varphi$. If also a variable not in $A$ is set, then UCP can be characterized by $G^{d}(\varphi,A)$ as follows.
Lemma 5.3.
Assume $\varphi\subseteq\Psi_{n,k}$, a $(k-1)$-set $A\subseteq X$, and let $\alpha=\bigwedge_{x\in A}x$. Consider a partial assignment $\beta\in\{\neg x_{\ell},x_{\ell}\}$ setting a variable $x_{\ell}\in X\setminus A$ to 00 (the negative literal) or to $1$ (the positive literal). Then $\mathrm{UCP}(\varphi\wedge\alpha\wedge\beta)$ is consistent and we have
  • $\mathrm{UCP}(\varphi\wedge\alpha\wedge\neg x_{\ell})$ extends $\alpha$ precisely by setting to 00 the variables $x_{i}$ such that $G^{d}(\varphi,A)$ contains a path from $x_{i}$ to $x_{\ell}$.
  • $\mathrm{UCP}(\varphi\wedge\alpha\wedge x_{\ell})$ extends $\alpha$ at least by setting to $1$ the variables $x_{i}$ such that $G^{d}(\varphi,A)$ contains a path from $x_{\ell}$ to $x_{i}$.
Proof.
By (12), setting variables in $A$ to $1$ and all variables in $X\setminus A$ to 00 or all of them to $1$ is consistent with $f_{n,k}$. Hence, it is consistent also with the formula $\Psi_{n,k}$ representing $f_{n,k}$ and its subset $\varphi$. It follows that $\mathrm{UCP}(\varphi\wedge\alpha\wedge\beta)$ is consistent.
When simplifying a subset of $\Psi_{n,k}$ using $\alpha$, the clauses containing a positive literal on a variable from $A$ become satisfied and, hence, are removed. Since $|A|=k-1$ and every clause of $\varphi$ contains $k$ negative literals, the clauses remaining in the simplified formula contain one positive and at least one negative literal. In particular, this is a definite Horn formula without unit clauses. In such a formula, setting a variable $x_{\ell}\in X\setminus A$ to 00 can derive a value of a new variable by UCP only using a clause containing the positive literal $x_{\ell}$ and exactly one negative literal. This negative literal is the derived one. Since the derived value of the variable in this literal is 00, UCP can derive only negative literals even if the propagation is iterated.
Simplifying a clause $C\in\varphi$ by $\alpha$ produces a clause with one negative literal only if $C=\mathrm{neg}(A)\cup\{\neg x_{i},x_{j}\}$ for some $i,j$ and the resulting clause is $\{\neg x_{i},x_{j}\}$. Hence, the clauses of the formula $\varphi$ simplified by $\alpha$ which are relevant for propagating assignments of variables in $X\setminus A$ to 00 form the 2-CNF formula
$\varphi_{A}=\{\{\neg x_{i},x_{j}\}\mid(x_{i},x_{j})\in G^{d}(\varphi,A)\}\;.$
If $x_{j}=0$, then UCP on $\varphi_{A}$ derives $x_{i}=0$ in one step if and only if $\{\neg x_{i},x_{j}\}\in\varphi_{A}$ and this is equivalent to $(x_{i},x_{j})\in G^{d}(\varphi,A)$. Hence, propagation of zeros in $\varphi_{A}$ is precisely represented by the reverse reachability in $G^{d}(\varphi,A)$ which implies the statement concerning $\mathrm{UCP}(\varphi\wedge\alpha\wedge\neg x_{\ell})$.
The statement concerning $\mathrm{UCP}(\varphi\wedge\alpha\wedge x_{\ell})$ is clear, since $\varphi_{A}$ is a subset of the formula $\varphi$ simplified by $\alpha$ and the clauses of $\varphi_{A}$ allow propagation of values $1$ along all directed paths in $G^{d}(\varphi,A)$. ∎
Proposition 5.4.
A formula $\varphi\subseteq\Psi_{n,k}$ is ucp-equivalent to $\Psi_{n,k}$ if and only if for every $(k-1)$-set $A\subseteq X$ the graph $G^{d}(\varphi,A)$ is a strongly connected graph on $X\setminus A$.
Proof.
Assume, $\varphi\subseteq\Psi_{n,k}$ is ucp-equivalent to $\Psi_{n,k}$. Let $A$ be any $(k-1)$-subset of $X$ and let $\alpha=\bigwedge_{x\in A}x$. By Lemma 5.3, reachability of the vertices in the graphs $G^{d}(\Psi_{n,k},A)$ and $G^{d}(\varphi,A)$ represent propagating assignments of the variables in $X\setminus A$ to 00 in the formulas $\Psi_{n,k}\wedge\alpha$ and $\varphi\wedge\alpha$, respectively. By assumption, these formulas are ucp-equivalent, so the two graphs have the same transitive closure. Since the graph $G^{d}(\Psi_{n,k},A)$ is the complete directed graph on $X\setminus A$, the graph $G^{d}(\varphi,A)$ is strongly connected.
Assume that $G^{d}(\varphi,A)$ is strongly connected. Let $C\in\Psi_{n,k}\setminus\varphi$ and let us prove that $C$ is absorbed by $\varphi$. Let $l\in C$ and consider a partial assignment $\gamma$ of the variables falsifying all the literals in $C\setminus\{l\}$. We have to show that UCP on $\varphi\wedge\gamma$ derives $l$.
If $l$ is positive, let $g\in C\setminus\{l\}$ be arbitrary. If $l$ is negative, let $g$ be the positive literal contained in $C$. For a suitable $(k-1)$-set of variables $A$, we have $C=\mathrm{neg}(A)\cup\{g,l\}$ and $g$ and $l$ have the opposite sign. Since $\gamma$ falsifies all the literals in $\mathrm{neg}(A)$ and $g$, we have $\gamma=\alpha\wedge\neg g$ where $\alpha=\bigwedge_{x\in A}x$.
Assume that the variable in $g$ is $x_{i}$ and the variable in $l$ is $x_{j}$. Hence, we have either $g=x_{i}$ and $l=\neg x_{j}$ or $g=\neg x_{i}$ and $l=x_{j}$. By assumption, $G^{d}(\varphi,A)$ contains paths between $x_{i}$ and $x_{j}$ in both directions. If $g=x_{i}$, then $\gamma=\alpha\wedge\neg x_{i}$ and Lemma 5.3 implies that UCP on $\varphi\wedge\gamma$ derives $\neg x_{j}=l$. If $g=\neg x_{i}$, then $\gamma=\alpha\wedge x_{i}$ and Lemma 5.3 implies that UCP on $\varphi\wedge\gamma$ derives $x_{j}=l$.
We obtained that every clause $C\in\Psi_{n,k}\setminus\varphi$ is absorbed by $\varphi$. It follows that $\Psi_{n,k}$ is absorbed by $\varphi$. Since $\varphi$ is a subset of $\Psi_{n,k}$, it is ucp-equivalent to $\Psi_{n,k}$. ∎
Proposition 5.5.
The size (number of clauses) of an arbitrary CNF formula ucp-equivalent to $\Psi_{n,k}$ is at least $n\,C(n-1,k,k-1)$.
Proof.
Let $\varphi^{*}$ be a smallest CNF formula ucp-equivalent to $\Psi_{n,k}$. By Lemma 5.1, without loss of generality, we can assume that $\varphi^{*}$ consists only of its prime implicates and, hence, is a subset of $\Psi_{n,k}$. For every variable $x\in X$, let
$\mathcal{S}_{x}=\{B\mid\mathrm{neg}(B)\vee x\in\varphi^{*}\}\;.$
Clearly, $\mathcal{S}_{x}$ is a system of $k$-sets $B\subseteq X\setminus\{x\}$.
Let $x\in X$ and let $A$ be an arbitrary $(k-1)$-subset of $X\setminus\{x\}$. By Proposition 5.4, vertex $x$ has non-zero in-degree in the graph $G^{d}(\varphi^{*},A)$, so there is a vertex $y\in X\setminus A$ different from $x$ such that $(y,x)$ is an edge of the graph. By construction of $G^{d}(\varphi^{*},A)$, $\mathrm{neg}(A)\cup\{\neg y,x\}$ is a clause of $\varphi^{*}$. The set $B=A\cup\{y\}$ is a member of $\mathcal{S}_{x}$ and $A\subseteq B$. It follows that $\mathcal{S}_{x}$ is a system of $k$-subsets of an $(n-1)$-set covering all $(k-1)$-subsets. Hence, $|\mathcal{S}_{x}|\geq C(n-1,k,k-1)$ and, since $|\varphi^{*}|=\sum_{x\in X}|\mathcal{S}_{x}|$, the statement follows. ∎
Let us recall that both the formulas $\varphi^{\ell}$ and $\varphi^{*}$ needed for the main result will be constructed as hypergraph formulas. Proposition 5.4 characterizes general subsets of $\Psi_{n,k}$ which are ucp-equivalent to the whole formula using the directed graphs $G^{d}(\varphi,A)$. It appears that for hypergraph formulas $\theta(H)\subseteq\Psi_{n,k}$ the following simpler condition using an undirected graph is sufficient for the same purpose.
Definition 5.6.
We say that a $(k+1)$-uniform hypergraph $H$ on the set of vertices $X$ has connected restrictions, if for every $(k-1)$-set $A\subseteq X$, the graph
$G(H,A)=\{e\setminus A\mid A\subseteq e\in H\}$ (13)
is a connected graph on the set $X\setminus A$.
Proposition 5.7.
If $H$ is a $(k+1)$-uniform hypergraph on $X$, then $\theta(H)$ is ucp-equivalent to $\Psi_{n,k}$ if and only if $H$ has connected restrictions.
Proof.
Since $\theta(H)$ is a subset of $\Psi_{n,k}$, we can use the characterization of formulas ucp-equivalent to $\Psi_{n,k}$ from Proposition 5.4.
Let $A$ be an arbitrary $(k-1)$-subset of $X$. By construction of $\theta(H)$, if $G^{d}(\theta(H),A)$ contains the directed edge $(x_{i},x_{j})$, then it also contains the directed edge $(x_{j},x_{i})$. Moreover, these two edges are contained in $G^{d}(\theta(H),A)$ if and only if $G(H,A)$ contains the undirected edge $\{x_{i},x_{j}\}$. Hence, the graph $G^{d}(\theta(H),A)$ is strongly connected if and only if $G(H,A)$ is connected. Since this is satisfied for arbitrary $(k-1)$-subset $A\subseteq X$, the statement follows. ∎
Note that $\Psi_{n,k}$ is a hypergraph formula, since it is $\theta(H_{0})$, where $H_{0}$ is the complete $(k+1)$-uniform hypergraph on $X$. The ucp-irredundant formula $\varphi^{\ell}$ can be obtained from an easy to define subset of $H_{0}$.
Definition 5.8.
Let $H_{1}$ consist of all $(k+1)$-subsets of $X$ containing the variable $x_{1}$ and let $\varphi^{\ell}=\theta(H_{1})$.
The proportion of the $(k+1)$-subsets of $X$ containing the variable $x_{1}$ in the set of all these subsets is $(k+1)/n$. For the main result, we assume $k=n/2+O(1)$, so the size of $\varphi^{\ell}$ is roughly one half of the size of $\Psi_{n,k}$.
Proposition 5.9.
Formula $\varphi^{\ell}$ is ucp-irredundant and ucp-equivalent to $\Psi_{n,k}$.
Proof.
The hypergraph $H_{1}$ has connected restrictions, since for every $(k-1)$-set $A$ not containing $x_{1}$, the graph $G(H_{1},A)$ is a star on $X\setminus A$ whose central vertex is $x_{1}$. If $x_{1}\in A$, then $G(H_{1},A)$ is a complete graph on $X\setminus A$. It follows that $\varphi^{\ell}$ is ucp-equivalent to $\Psi_{n,k}$ by Proposition 5.7.
Let us prove that $\varphi^{\ell}$ is ucp-irredundant. For this purpose, let $C\in\varphi^{\ell}$ be arbitrary and let us prove that $\varphi^{\ell}\setminus\{C\}$ is not ucp-equivalent to $\Psi_{n,k}$. Let $V$ be the set of variables used in $C$ except of $x_{1}$, so we have $|V|=k$. Let $\gamma$ be the partial assignment of the variables in $V$ such that all the literals on variables from $V$ in $C$ are falsified. UCP on the clause $C$ together with $\gamma$ derives a value of $x_{1}$ which makes the literal on $x_{1}$ in $C$ satisfied. Since $C$ is an implicate of $\Psi_{n,k}$, every formula ucp-equivalent to $\Psi_{n,k}$ should derive the same value of $x_{1}$ as $C$. The proof will be finished by proving that $\varphi^{\ell}\setminus\{C\}$ does not satisfy this.
Depending on whether $C$ contains the literal $x_{1}$ or $\neg x_{1}$, the partial assignment $\gamma$ assigns all variables in $V$ the value $1$ or all variables in $V$ are assigned the value $1$ except of one which is assigned 00. In both cases, the following is true. Every clause $D$ in $\varphi^{\ell}\setminus\{C\}$ on the same set of variables as $C$ is satisfied by $\gamma$. The reason is that two of the variables in $C$ have different sign in $C$ and $D$. At least one of them is not $x_{1}$. The literal with this variable in $C$ is falsified, so its negation in $D$ is satisfied. Every clause in $\varphi^{\ell}\setminus\{C\}$ on a set of variables different from the set of variables of $C$ contains at least two variables not assigned by $\gamma$, namely $x_{1}$ and some variable not in $V\cup\{x_{1}\}$. Such a clause cannot be used for UCP. Altogether, UCP on $\varphi^{\ell}\setminus\{C\}$ with the assignment $\gamma$ does not derive a value of any new variable, in particular, a value of $x_{1}$. ∎
Construction of a hypergraph $H$ such that $\varphi^{*}=\theta(H)$ is a sufficiently small formula ucp-equivalent to $\Psi_{n,k}$ is more complicated and we use results proven in the next two sections for this purpose.
6 Union of random permutations of a graph
In the next section we prove the existence of a sufficiently small $(k+1)$-uniform hypergraph on the set $X$ with connected restrictions, so that Proposition 5.7 then allows to prove existence of the required formula $\varphi^{*}$. Using known results on covering systems, it is possible to construct a hypergraph $H$ satisfying a weaker condition, namely, that the graph $G(H,A)$ has no isolated vertices for every $A\subseteq X$ of size $k-1$. A possible way how to extend a graph with no isolated vertices to a connected graph is to take a union of a constant number of its copies obtained by a random permutation of the vertices. The purpose of this section is to prove a small upper bound on the probability that this union fails to be connected. If this probability is small enough, then the process can be used in parallel for almost all sets $A$ which we consider and allows to prove the existence of the required hypergraph.
The following definition will be used in the next section for hypergraphs and this section uses its special case for graphs which are $2$-uniform hypergraphs.
Definition 6.1.
Let $H$ be a hypergraph on a set of vertices $V$. Consider a permutation $\pi:V\to V$. We denote by $\pi(H)$ the hypergraph on the set of vertices $V$ such that for every set $e\subseteq V$, $\pi(e)$ is a hyperedge of $\pi(H)$ if and only if $e$ is a hyperedge of $H$.
Intuitively, graph or hypergraph $\pi(H)$ is obtained by moving each vertex $u$ of $H$ to $\pi(u)$ and moving the edges correspondingly. By a random permutation of $H$, we mean $\pi(H)$, where $\pi$ is chosen at random from the uniform distribution on permutations of $V$.
An undirected graph $G=(V,E)$ is connected if and only if for every partition of $V$ into non-empty sets $Z$ and $V\setminus Z$, there is an edge connecting a vertex in $Z$ and a vertex in $V\setminus Z$. We need an upper bound on the probability of the complement of this event for a fixed set $Z\subseteq V$ and a graph obtained as a random permutation of a given graph. Equivalently, we can consider a fixed graph and a random set $Z$ of a given size. Clearly, there is no edge between $Z$ and $V\setminus Z$ if and only if $Z$ is a union of the connected components of the graph.
Assume a graph on the set of vertices $V$ of size $m$ with no isolated vertices and with $r$ connected components of sizes $a_{i}\geq 2$, $i=1,\ldots,r$. The following lemma provides an upper bound on the number of subsets $Z$ of $V$ of size $d$ which are unions of connected components of the graph. In the simple case when $m$ and $d$ are even and $a_{i}=2$ for $i=1,\ldots,m/2$, the number of these sets is exactly (${m/2\choose d/2}$. In the general case, a similar estimate can be obtained. The considered subsets of $V$ are encoded by the sets $I$ satisfying (14). Since the sets $I$ are pairwise incomparable we can use LYM inequality [19] implying an upper bound on the number of sets of different sizes in a collection of pairwise incomparable sets.
Lemma 6.2 ([20]).
Assume integers $a_{i}\geq 2$ for $i=1,\ldots,r$ and let $m=\sum_{i=1}^{r}a_{i}$. Then for every $2\leq d\leq\lfloor m/2\rfloor$, the number of sets $I\subseteq\{1,\ldots,r\}$, such that
$\sum_{i\in I}a_{i}=d$ (14)
is at most
(${\lfloor m/2\rfloor\choose\lfloor d/2\rfloor}\;.$
Proof.
The sets $I$ satisfying (14) are pairwise incomparable and have size at most $\lfloor d/2\rfloor$. Let $c_{j}$ for $j=1,\ldots,\lfloor d/2\rfloor$ be the number of these sets of size $j$. By LYM inequality [19], we have
$\sum_{j=1}^{\lfloor d/2\rfloor}\frac{c_{j}}{{r\choose j}}\leq 1\;.$
Since $r\leq\lfloor m/2\rfloor$, we have also
$\sum_{j=1}^{\lfloor d/2\rfloor}\frac{c_{j}}{{\lfloor m/2\rfloor\choose j}}\leq 1$
and, since $\lfloor d/2\rfloor\leq\lfloor m/2\rfloor/2$, this implies that the number of sets $I$ satisfying (14) is at most
$\sum_{j=1}^{\lfloor d/2\rfloor}c_{j}\leq{\lfloor m/2\rfloor\choose\lfloor d/2\rfloor}$
as required. ∎
Lemma 6.3.
Let $G=(V,E)$ be an undirected graph with no isolated vertices where $|V|=m$, let $22\leq d\leq m/$ 2, and let $Z$ be a subset of $V$ of size $d$. The probability that a random permutation $\pi(G)$ of $G$ has no edge connecting $Z$ and $V\setminus Z$ is at most
(${\lfloor m/2\rfloor\choose\lfloor d/2\rfloor}{m\choose d}^{-1}\;.$
Proof.
Consider a random permutation $\pi$ of the vertices of $G$. Graph $\pi(G)$ has no edge between $Z$ and $V\setminus Z$ if and only if the preimage $\pi^{-1}(Z)$ is a union of connected components of $G$. The preimages $\pi^{-1}(Z)$ are uniformly distributed over subsets of $V$ of size $d$. The number of subsets of $V$ of size $d$ that are a union of connected components in $G$ is equal to the number of solutions of the equation (14) where $r$ is the number of connected components of $G$ and $a_{i}$ are their sizes. Since $G$ has no isolated vertices, we have $a_{i}\geq 2$ for $i=1,\ldots,r$ and, hence, the statement follows from Lemma 6.2. ∎
Proposition 6.4.
Let $G_{1},\ldots,G_{s}$, $s\geq 3$, be graphs on the same set of $m$ vertices each of which has no isolated vertices. If $s$ is fixed and $m$ tends to infinity, then the probability that the union of independent random permutations of $G_{1},\ldots,G_{s}$ is not a connected graph is at most $O(m^{2-s})$.
Proof.
Let $V$ be the common set of vertices of the graphs. The probability of the event that the union of independent permutations of $G_{1},\ldots,G_{s}$ is not a connected graph is at most the sum of the probabilities over subsets $Z\subseteq V$ of size between $2$ and $m/2$ of the event that the union contains no edge between $Z$ and $V\setminus Z$. Using $d=|Z|$ as a summation index and using Lemma 6.3, this is at most
$\sum_{d=2}^{\lfloor m/2\rfloor}p_{d}$
where
$p_{d}={\lfloor m/2\rfloor\choose\lfloor d/2\rfloor}^{s}{m\choose d}^{1-s}\;.$ (15)
For a few smallest values of $d$, the following simple estimate is sufficient. If $d$ and $s$ are fixed and $m\to\infty$, we have
$p_{d}=O\left(m^{sd/2+(1-s)d}\right)=O\left(m^{(2-s)d/2}\right)\;.$ (16)
For larger values of $d$, we estimate $p_{d}$ using a more precise bound on binomial coefficients. Denote
$b(t,d)=\left(\frac{t}{d}\right)^{d}\left(\frac{t}{t-d}\right)^{t-d}=\mathrm{e}^{t\mathrm{H}(d/t)}$
where $\mathrm{H}(x)=-x\ln x-(1-x)\ln(1-x)$ is the entropy function. Since $\mathrm{H}(x)$ is increasing for $20\leq x\leq 1/$ 2, $b(t,d)$ is increasing in $d$ if $t$ is fixed and $21\leq d\leq t/$ 2. Since $\mathrm{H}(x)$ is strictly concave and $\mathrm{H}(0)=0$, $\mathrm{H}(x)/x$ is decreasing for $0<x\leq 1$. By substitution $x=d/t$, we obtain that $b(t,d)$ is increasing in $t$ if $d$ is fixed. By a well-known upper bound on binomial coefficients and using $d\leq m/2$, we get
(${\lfloor m/2\rfloor\choose\lfloor d/2\rfloor}\leq b(\lfloor m/2\rfloor,\lfloor d/2\rfloor)\leq b(m/2,d/2)=b(m,d)^{1/2}\;.$
By a well-known lower bound on binomial coefficients, we have
$\frac{1}{m+1}b(m,d)\leq{m\choose d}\;.$
Using these bounds and (15), we obtain
$p_{d}\leq(m+1)^{s-1}b(m,d)^{1-s/2}\;.$ (17)
Let $d_{0}=\lceil 4(s-1)/(s-2)\rceil$. Using (16) for $d\geq 2$, we obtain
$\sum_{d=2}^{d_{0}-1}p_{d}=O(m^{2-s})\;.$ (18)
If $d_{0}\leq d\leq m/2$, then $b(m,d)\geq b(m,d_{0})\geq(m/d_{0})^{d_{0}}$ and by (17)
$p_{d}\leq(m+1)^{s-1}(m/d_{0})^{d_{0}(1-s/2)}=O(m^{1-s})\;.$
Together with (18), we obtain
$\sum_{d=2}^{\lfloor m/2\rfloor}p_{d}=O(m^{2-s})$ (19)
and the proof is finished. ∎
7 Size of a hypergraph with connected restrictions
The formula $\varphi^{*}$ needed for the main result will be constructed as a hypergraph formula. By Proposition 5.7, in order to obtain the required formula, we need a sufficiently small hypergraph with connected restrictions. In this section, we prove the existence of such a hypergraph whose size is only by a constant factor larger than the size of a suitable covering system.
We start from a covering system $H$ and show that as a hypergraph it satisfies a condition related to but weaker than having connected restrictions. As suggested in the previous section, we consider several independent random permutations of its vertices. With high probability, the union $H^{*}$ of the permuted hypergraphs has almost all restrictions connected or, more formally, $G(H^{*},A)$ is connected for almost all sets $A$ of size $k-1$. The final hypergraph $H^{**}$ is then obtained by extending $H^{*}$ with a limited number of hyperedges so that it has connected restrictions and the required upper bound on its size is achieved.
Proposition 7.1.
For every large enough $n$ and an integer $k=n/2+O(1)$, there is a $(k+1)$-uniform hypergraph on $n$ vertices of size at most $6\,C(n,k+1,k)$ with connected restrictions.
Proof.
Let $|X|=n$ and let $H$ be a $(k+1)$-uniform hypergraph of size $C(n,k+1,k)$ covering all $k$-subsets of $X$. Let $s=5$, let $\pi_{1},\ldots,\pi_{s}$ be independent random permutations of $X$ chosen from the uniform distribution, and let $H^{*}=\pi_{1}(H)\cup\ldots\cup\pi_{s}(H)$.
If $A$ is a $(k-1)$-subset of $X$ and $x\in X\setminus A$, then by assumption, there is a hyperedge $B\in H$ containing $A\cup\{x\}$, so $B=A\cup\{x,y\}$ for some $y\in X\setminus A$. In particular, vertex $x$ has non-zero degree in $G(H,A)$. This argument applies also to each of the graphs $G(\pi_{i}(H),A)$, so each of them has no isolated vertices. The graph $G(H^{*},A)$ is a union of $s$ graphs obtained in this way, however, we cannot directly use Proposition 6.4. The graph $G(\pi_{i}(H),A)$ is isomorphic to $G(H,\pi_{i}^{-1}(A))$ and, hence, it is not, in general, a random permutation of a fixed graph on $X\setminus A$. In order to overcome this, we partition the probability space of all $s$-tuples of permutations of $X$ by a set of mutually exclusive conditions such that the assumption of Proposition 6.4 is satisfied by the conditional distribution of the $s$-tuple of graphs $G(\pi_{i}(H),A)$, $i=1,\ldots,s$ under any of these conditions. Namely, for an arbitrary $s$-tuple $(A_{1}^{\prime},\ldots,A_{s}^{\prime})$ of $(k-1)$-subsets of $X$ let $c(A_{1}^{\prime},\ldots,A_{s}^{\prime})$ be the condition $\pi_{i}(A_{i}^{\prime})=A$ for all $i=1,\ldots,s$. Clearly, these conditions represent a partition of the probability space and, since $\pi_{i}$ are independent and uniformly distributed, all of these conditions have the same probability.
Fix an arbitrary $s$-tuple $(A_{1}^{\prime},\ldots,A_{s}^{\prime})$ and consider the conditional distribution under $c(A_{1}^{\prime},\ldots,A_{s}^{\prime})$. For a fixed $i$ and all permutations $\pi_{i}$ satisfying $\pi_{i}(A_{i}^{\prime})=A$, the graphs $G(\pi_{i}(H),A)$ are isomorphic, since all of them are isomorphic to $G(H,A_{i}^{\prime})$. Choose any of the graphs $G(\pi_{i}(H),A)$ as $G_{i}$. The distribution of $G(\pi_{i}(H),A)$ is the same as the distribution of random permutations of $G_{i}$. Moreover, these random permutations are independent for $i=1,\ldots,s$. Hence, we can use Proposition 6.4 with $V=X\setminus A$ to obtain that $G(H^{*},A)$ is not connected with probability at most $O((n-k+1)^{2-s})$ under any of the considered conditions and, hence, also unconditionally.
A $(k-1)$-set $A$ is called “bad”, if the graph $G(H^{*},A)$ is not connected and let $t$ be the random variable equal to the number of such sets. By the argument above, we have
$\mathbf{E}\;t=O\left((n-k+1)^{2-s}\right){n\choose k-1}\;.$ (20)
For each of the $t$ “bad” sets $A$, we add several $(k+1)$-sets to $H^{*}$ in order to add edges to $G(H^{*},A)$ so that this graph becomes connected. The number of $(k+1)$-sets needed for this is at most the number of the connected components of $G(H^{*},A)$ minus $1$ which is at most $(n-k-1)/2$. Hence, the resulting $(k+1)$-uniform hypergraph $H^{**}$ has size at most $s|H|+t(n-k-1)/2$. By (20), there is a choice of permutations $\pi_{i}$ for $i=1,\ldots,s$, such that
$t=O\left((n-k+1)^{2-s}\right){n\choose k-1}=O\left((n-k+1)^{2-s}\right){n\choose k}$
Since $n-k=\Theta(k)$, we can conclude
$|H^{**}|\leq s|H|+O\left(k^{3-s}\right){n\choose k}\;.$
Using (1) and $s=5$, we get
$|H^{**}|\leq(s+1)|H|\leq 6\,C(n,k+1,k)$
if $n$ and, hence, also $k$ are sufficiently large. This finishes the proof, since $H^{**}$ has connected restrictions. ∎
8 Proof of the results
In this section, we summarize the results of the previous sections. First, we finish the proof of Proposition 4.3 on a lower and an upper bound on the minimum size of a formula ucp-equivalent to $\Psi_{n,k}$.
Proof of Proposition 4.3.
The lower bound is Proposition 5.5. For the upper bound, assume a large enough $n$ and $k=n/2+O(1)$. Our goal is to prove that there is a formula ucp-equivalent to $\Psi_{n,k}$ and of size at most $6(k+1)C(n,k+1,k)$. For this purpose, we use a hypergraph formula. Consider a hypergraph $H$ of size at most $6\,C(n,k+1,k)$ with connected restrictions guaranteed by Proposition 7.1. By Proposition 5.7, $\theta(H)$ is ucp-equivalent to $\Psi_{n,k}$ and its size is $(k+1)|H|$ as required. Using (2) and the bound (5) or (7), we obtain the required asymptotic estimate. ∎
The constructions presented up to now imply Theorem 4.1 as follows.
Proof of Theorem 4.1.
Let $\varphi^{\ell}$ be the formula from Definition 5.8. This formula is ucp-irredundant and ucp-equivalent to $\Psi_{n,k}$ by Proposition 5.9. Assuming $k=n/2+O(1)$, its size is
$|\varphi^{\ell}|=(k+1){n-1\choose k}=(k+1)\left(1-\frac{k}{n}\right){n\choose k}=\Theta(n){n\choose k}\;.$
Let $\varphi^{*}\subseteq\Psi_{n,k}$ be a formula ucp-equivalent to $\Psi_{n,k}$ of size satisfying the upper bound from Proposition 4.3. The statement follows by comparing the size of the formulas $|\varphi^{\ell}|$ and $|\varphi^{*}|$. ∎
9 Conclusion and further research
The paper introduces ucp-equivalence on CNF formulas and investigates the ratio of the size of a ucp-irredundant formula and the size of a smallest ucp-equivalent formula. The paper shows that a general upper bound on this ratio is at most $n^{2}$ (Proposition 3.5) and at least $\Omega(n/\ln n)$ (Theorem 4.1). The gap between these bounds is quite large. An interesting open problem is to reduce this gap. It is plausible to assume that the upper bound can be improved. Another interesting question is whether the lower bound $\Omega(n/\ln n)$ or an even better bound can be proven using a different type of functions and without the use of probabilistic method.
Acknowledgement. The author would like to thank Diana Piguet for a fruitful discussion on using hypergraphs, in particular, for suggesting Definition 4.4 which later appeared to be used also for other purposes, and Matas Šileikis for suggesting Lemma 6.2 which improves a slightly weaker bound on the same quantity by the author. The author acknowledges institutional support RVO:67985807.